### v1

# DBudgetKV: Dynamic Budget in KV Cache Compression for Ensuring Optimal Performance

Abstract
To alleviate memory burden during inference of large language models (LLMs), numerous studies have focused on compressing the KV cache by exploring aspects such as attention sparsity. However, these techniques often require a pre-defined cache budget; as the optimal budget varies with different input lengths and task types, it limits their practical deployment accepting open-domain instructions. To address this limitation, we propose a new KV cache compression objective: to always ensure the full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible. To achieve this goal, we introduce a novel KV cache compression method dubbed DBudgetKV, which features an attention-based metric to signal when the remaining KV cache is unlikely to match the full-cache performance, then halting the pruning process. Empirical evaluation spanning diverse context lengths, task types, and model sizes suggests that our method achieves lossless KV pruning effectively and robustly, exceeding 25% compression ratio on average. Furthermore, our method is easy to integrate within LLM inference, not only optimizing memory space, but also showing reduced inference time compared to existing methods.
DBudgetKV: Dynamic Budget in KV Cache Compression
for Ensuring Optimal Performance
Xuanfan Nithanks: Equal contributions. Partial work was done during Xuanfan’s internship at WeChat AI.   Liyan Xu11footnotemark: 133footnotemark: 3   Chenyang Lyu  Longyue Wang Mo Yu33footnotemark: 3   Lemao Liu33footnotemark: 3   Fandong Meng33footnotemark: 3   Jie Zhou33footnotemark: 3   Piji Lithanks: Corresponding author. 33footnotemark: 3Pattern Recognition Center, WeChat AI xuanfanni@gmail.comliyanlxu@tencent.comlipiji.pz@gmail.com
1 Introduction
As most large language models (LLMs) follows autoregressive generation Minaee et al. (2024); Chang et al. (2023); OpenAI (2023), they rely on KV Cache to store intermediate states during inference. Specifically, LLMs access the cached key and value vectors of past tokens during the attention calculation, thereby speeding up inference by re-using these KV cache Vaswani et al. (2017); Ainslie et al. (2023); Shazeer (2019). However, as the model size and the input length increase, the memory required to maintain the KV cache also grows proportionally. Since these vectors are usually stored within GPU memory, managing the KV cache efficiently has become crucial to mitigate the overall memory consumption and computational overhead during LLM inference. For instance, Llama3 8B model Dubey et al. (2024) requires 1 GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
Consequently, recent works have proposed to effectively reduce KV cache through pruning, leveraging the sparsity of the attention mechanism: certain positions are more pivotal during the generation process, while those less important ones could be pruned with minimal performance degradation; thus, the full KV cache is not always necessary. These observations have led to several optimization methods, such as H2O Zhang et al. (2023), ScissorHands Liu et al. (2023), StreamingLLM Xiao et al. (2024), and FastGen Ge et al. (2024), where they aim to identify and retain only the most salient token positions, discarding less critical ones based on certain pruning criteria. Alongside, other orthogonal paradigms have also been proposed without hard-pruning, such as KVMerger Wang et al. (2024) and D2O Wan et al. (2024) that merge KV vectors, and MLA DeepSeek-AI et al. (2024) that operates attention in the latent space.
While previous pruning methods have successfully demonstrated that a decent size of KV cache could be discarded in practice, there exists one common limitation that has not been addressed yet: these pruning techniques typically require a pre-defined KV cache budget, of which its optimal threshold could vary significantly according to specific tasks or inputs. Hence, in real-world scenarios, one would have to carefully tune many budget thresholds for diverse domains, or set a uniform threshold but likely suffering large degradation on certain tasks. We reckon that neither choice is a satisfactory solution, hindering the widespread adoptation of KV pruning techniques.
To further illustrate, Table 1 shows that when the budget is set to 50%, Llama3-8B-Instruct using H2O achieves over 98% of the full-cache performance on NarrativeQA Kociský et al. (2018), whereas under the same conditions on GSM8K Cobbe et al. (2021), a widely-used math benchmark, the performance drastically falls to less than 42% of the full-cache performance.
To address the aforementioned limitation, in this work, we introduce a new objective for KV cache compression: our method aims to strike towards full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible. Towards this objective, we propose an adaptive pruning method, dubbed DBudgetKV, which Dynamically adjusts KV cache retention per input, without the need of pre-defined memory Budget. Particularly, DBudgetKV automatically identifies input-dependent KV cache pruning that leads to minimal performance degradation. A key implication is that the method tends to realize a higher compression ratio for simpler tasks, while allocating more cache resources for complicated tasks.
Our method features a two-step process: 1) after the prefilling stage, rank all input tokens according to an importance metric; 2) discard the KV cache of each token sequentially, until hitting a stopping criterion. In choosing the importance metric and the stopping criterion, DBudgetKV takes into account two important factors from previous works: i) position bias, where tokens at the beginning and the end are generally more significant than those in the middle Xiao et al. (2024); Jiang et al. (2024b); ii) attention score, where tokens receiving higher scores usually contribute more information to subsequent generation Minaee et al. (2024); Chang et al. (2023); Zhang et al. (2023).
We then propose DBudgetKV of empirical efficacy and low overhead. Concretely, tokens are ranked solely by their positions, with no additional computation involved. The ranking strategy signals high significance to the beginning and the end positions, where the initial tokens are the most important, followed by the remaining tokens in a reversing order. Next in the pruning process, we design an attetion-based metric to determine the stopping condition, such that the pruning halts when this metric signals that the remaining KV cache is unlikely to match the full-cache capacity, thereby fulfilling our objective to approximate full performance with dynamic cache pruning.
More specifically, our attention metric compares the difference between the Frobenius form of the full attention matrix and the reduced matrix by setting pruned positions’ score to zero. When the difference exceeds a universal threshold, the process is halted. Especially, this metric empirically correlates well with the generation performance regardless of inputs, serving as an estimator to bridge the pruning state and the subsequent generation. Furthermore, the entire process is implemented efficiently through PyTorch’s built-in operators, where we show that DBudgetKV not only optimizes memory space, but also acheives time reduction compared to the full-cache inference and three KV pruning methods.
Figure 1: The overall workflow of DBudgetKV in Section 3.2. Initially, tokens are ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix. The KV cache for the remaining tokens are then preserved.
Experimental results on 13 datasets varying diverse context lengths and tasks, e.g. mathematical and commonsense reasoning, reading comprehension and coding, demonstrate that DBudgetKV achieves our objective effectively and robustly. The resulting inference is on par or even surpasses the full-cache performance with multiple LLMs of different sizes, including Llama3 Dubey et al. (2024), Qwen2.5 Yang et al. (2024), and Mistral Jiang et al. (2024a). Meanwhile, the dynamic pruning of DBudgetKV allocates relatively high budget on math benchmarks but attains high compression ratio by up to 85% on comprehension tasks. The average budget size reaches 63.7% using Llama3-8B. The overall results suggest that DBudgetKV is able to accomplish lossless KV cache pruning with strong generalizability. Besides, it does not rely on specific model architectures and can coexist with other KV cache optimization techniques. We further conduct more experiments for in-depth analysis and ablation studies, depicting the rationality of our various design choices and hyperparameters.
Our contributions can be summarized below:
  • To the best of our knowledge, we are the first to propose the objective that ensures full-cache performance while maximizing the KV cache pruning, which holds greater practical values and deployment potentials.
  • We introduce DBudgetKV that employs a straightforward two-step process, where its dynamic pruning is designed with minimal degradation to the generation performance.
  • Extensive evaluation with diverse tasks and models indicates that compared to prior works, our method achieves lossless KV cache compression effectively, efficiently and robustly.
2 Related Work
KV Cache Compression
Recent advancements in KV cache compression have emerged to address memory constraints for model inference. Scissorhands Liu et al. (2023) observes the repetitive nature of attention patterns and the persistence of token importance. It then preserves the most important tokens based on attention scores to achieve KV cache compression. H2O Zhang et al. (2023) introduces dynamic eviction policies that strategically balance retention of recent tokens and historically significant "heavy hitters". StreamingLLM Xiao et al. (2024) enables infinite-length sequence processing without fine-tuning by stabilizing attention through landmark tokens. SnapKV Li et al. (2024) enhances compression efficiency through attention score-based selection and clustering of critical KV positions. FastGen Ge et al. (2024) proposes an adaptive KV cache management system that customizes retention strategies per attention head.
However, these methods require setting a fixed memory budget, which makes it difficult to consistently maintain full performance across different domains and datasets, limiting their practical application. In contrast, our method imposes no such constraints, targeting full-cache performance while allowing for dynamic budget size.
3 Methodology
We first elaborate our objective in Sec. 3.1 as a new direction for KV cache compression. Distinct from previous works, we aim to compress the KV cache as much as possible while ensuring LLMs’ performance remains intact. We then delineate our proposed approach along with its motivation behind in Sec. 3.2. Important implementation details are next discussed in Sec. 3.3.
3.1 Objective of KV Cache Compression
As mentioned in Sec. 1, most previous works on KV cache pruning require a pre-defined memory budget beforehand, either by fixing the number of cached positions directly Cai et al. (2024); Li et al. (2024), or retaining positions by a fixed ratio of input length Zhang et al. (2023); Wan et al. (2024). While these approaches are effective in certain circumstances, they may encounter practical challenges when deployed in real-world scenarios.
The biggest issue is that the optimal budgets for LLM inference in real-world scenarios are evidently infeasible to enumerate, as they could vary across tasks and domains, especially for open-domain instructions. For example, mathematical tasks are typically concise and require inference based on all given conditions, thus many positions may logically contribute to the generation, necessitating a relatively higher memory budget. Conversely, for article reading comprehension, LLM may only need a small set of critical positions for the answer generation, featuring a smaller budget in many cases.
As a side effect, inference with previous KV cache pruning is likely to experience performance instability across different inputs, which can be observed by our analysis in Sec. 4.2. Overall, the practical value is thereby diminished.
In this work, we resort to a dynamic paradigm that eliminates the need for manually setting a memory budget. We introduce a new objective for KV cache compression, aiming to automatically reduce the KV cache as much as possible, while fulfilling the condition that the method should always strike towards full-cache performance. In this way, one could utilize such pruning method off-the-shelf, without spending time and resources on tuning usage-specific hyperparameters, which we hope would further advance the research development of KV cache compression techniques.
3.2 DBudgetKV
Towards our objective, we design a novel ranking-based pruning method applied after the prefilling stage, consisting of two steps. First, all tokens are ranked according to an importance metric per Transformers layer. Then, the system keeps removing the KV cache sequentially, until hitting a stopping criterion. Ultimately, our proposed approach does not need to fix an optimal budget for storing KV cache. Rather, the stopping criterion is tied to the norm of attention matrices that works uniformly regardless of the inputs.
The motivation behind our two-step procedure comes from two phenomena brought by previous works. First, there exists position bias such that positions in the beginning and the end are typically important for the subsequent generation, as observed by StreamingLLM Xiao et al. (2024). Second, attention scores are usually an effective indicator of the token importance, as leveraged by many pruning techniques Liu et al. (2023); Zhang et al. (2023); Li et al. (2024). For our objective, we propose to combine the two important properties in our approach, described in details as follows.
Importance Ranking
Denote a LLM input sequence as $X=\{x_{1},x_{2},\dots,x_{n}\}$, where each Transformers layer originally consists of $n$ positions of KV vectors per attention head. As the pruning process starts from the least important KV vectors, all tokens are firstly ranked according to an importance strategy.
To this end, we resort to a simple ranking strategy leveraging the position bias, without undergoing any additional computation. As identified by StreamingLLM, a significant portion of attentions is allocated to the initial tokens, regardless of their relevance to the language modeling task, known as attention sinks. Inspired by this, we employ a purely position-based importance evaluation strategy: we assume that the first $m$ tokens are always crucial; while among the remaining $n-m$ tokens, those towards the end are more significant. By this strategy, the importance in $X$ is ranked from the highest to the lowest as $\{x_{1},x_{2},\dots,x_{m},x_{n},x_{n-1},\dots,x_{m+1}\}$, denoted as $\widehat{X}$, where $m$ is a chosen hyperparameter that works regardless of specific input sequences.
It should be noted that our choice on importance ranking is empirically supported rather than theoretically based. We have experimented other attention-based importance ranking, which are provided in Sec. 4.4. Our position-based strategy is shown both empirically effective and superior on computational overhead.
Stopping Criterion
With the initially ranked KV vectors, our method then sequantially decides whether to continue or terminate pruning, which needs an evaluation metric serving as the stopping criterion. Since we evidently cannot know the precise impact of a KV cache position before the generation, we require a metric that correlates well with the resulting performance change when discarding a position, serving as a bridge to ensure a minimal degradation upon the full KV-cache performance.
Inspired by Devoto et al. (2024), we utilize the Frobenius norm of the attention matrix that fits the requirement well empirically. For a matrix $M\in\mathbb{R}^{m\times n}$, the Frobenius norm of $M$ is the matrix-version of the L2 norm, expressed as $||M||_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}|M_{i,j}|^{2}}$.
We conduct preliminary experiments to validate its reliability. We employ Llama2-7B and randomly mask a certain portion of the KV cache, setting the corresponding values in the attention matrix $A\in\mathbb{R}^{n\times n}$ to zero. As illustrated in Figure 2, there is a certain degree of positive correlation between its Frobenius norm and the model’s performance, which are shown generalized across different domains.
Figure 2: Preliminary experiments on the relationship between the Frobenius norm difference (in percentage) obtained from randomly masking tokens, and the corresponding model performance on GSM8K.
As the input sequence length $n$ increases, the time and space overhead for norm calculation also grows significantly by $O(n^{2})$. Hence, we next aim to reduce the computational scale from the naive norm calculation. We observe that the Frobenius norm of the last row of $A$ also correlates with the model’s performance; equivalently saying, the attention distribution of the latest token is good enough to serve as an estimator. Formally, we utilize the last $k$ rows (latest $k$ tokens) of $A$ and reduce to a single attention vector $A^{\prime}\in\mathbb{R}^{1\times n}$. The score $s_{j}$ for a position $j\in[1,n]$ in $A^{\prime}$ is:
$s_{j}=\frac{\sum_{i=k}^{n}A_{i,j}}{\sum_{i=k}^{n}\mathbf{1}_{\{A_{i,j}\neq 0\}}}$ (1)
We then use the Frobenius norm of $A^{\prime}$, equivalently its L2 norm, as the evaluation metric, which reduces the overall complexity to $O(n)$.
Finally, for each position in the ranking order, our method compares the difference between the Frobenius norm of the original attention matrix $A$, and the Frobenius norm of the reduced attention vector $A^{\prime}$ of which the corresponding positions are set to 00. We set a universal threshold as 1%, such that when the new norm exceeds 1% difference from the original norm, it hits the stopping criterion. The KV vectors of the remaining positions are then all kept, while the preceding positions would have their KV cache discarded. In this way, our method achieves dynamic KV cache pruning while targeting to maintain the full-cache performance, assisted by the norm metric.
3.3 Implemetation Details
Though our proposed pruning process is conceptually sequential, it is implemented efficiently by PyTorch’s operators, such that the stopping positions of all layers are identified directly in parallel without any looping operations. Empirical analysis in Sec. 4.5 shows that our proposed method achieves higher efficiency compared to three popular baselines.
Specifically, we reorder $A^{\prime}$ based on Token Importance Ranking $\widehat{X}$ to obtain $A^{\prime}_{sort}$, where $A^{\prime}_{sort}[0]$ corresponds to the least important token and $A^{\prime}_{sort}[-1]$ corresponds to the most important token. Next, we compute the cumulative square-sum of each element in $A^{\prime}_{sort}$:
$A^{\prime}_{cumsum}[i]=\sqrt{\sum_{k=i}^{n}|A^{\prime}_{sort}[k]|^{2}}$ (2)
such that $A^{\prime}_{cumsum}[i]$ represents the Frobenius Norm of the attention matrix after removing all tokens to the left of $i^{th}$ token. We then divide each element in $A^{\prime}_{cumsum}$ by $||A^{\prime}||_{F}$ and subtract this value from 1 to determine the difference between each new matrix and $A^{\prime}$. The torch where operation allows us to directly identify the position of the token that satisfies the condition of not exceeding a 1% difference, ultimately yielding the set of tokens for which the KV cache is called to retain. The final kv cache compression algorithm of DBudgetKV is presented in Algorithm 1. Since PyTorch allows specifying the dimension for operations, DBudgetKV can run in parallel across multiple attention heads.
Retaining Bottom Layers
Additionally, we have discovered that applying DBudgetKV to the model’s 0th and 1st layers results in poor performance when the model operates under a high budget. We hypothesize that this may be due to the relatively uniform distribution of attention scores in these initial layers, leading to the premature discarding of tokens that are crucial for subsequent layers, ultimately causing the entire model’s inference to collapse. A more detailed analysis of this issue is provided in Appendix B. Based on this finding, we initiate the KV cache compression operations starting from the model’s 2nd layer.
4 Experiments
4.1 Experimental Settings
  Methods Math & Science CR Single-Doc QA Multi-Doc QA Summarization FSL Code   Avg.
GSM8K
GPQA
TheoQA
ThQA
CoQA
NrtvQA
Qasper
2WkMQA
Musique
QMSum
M-News
TriviaQA
Lcc
Llama3-8B-Instruct Full 75.28 29.02 21.29 25.59 52.74 24.06 43.91 35.33 14.77 22.27 27.37 70.26 19.16 100%
Oursk=1 76.50 30.13 23.03 26.09 52.86 23.44 37.38 36.21 15.96 21.95 27.88 64.24 19.31 +0.12%
Budget 93.2% 86.7% 92.8% 78.7% 81.5% 48.7% 46.4% 45.8% 43.7% 15.0% 76.4% 41.0% 78.0% 63.68%
H2O0.9 74.60 28.13 21.69 25.41 52.83 24.55 41.73 33.95 15.23 22.51 27.64 69.77 19.13 -0.39%
SLM0.9 72.78 28.79 20.75 25.15 52.83 23.75 43.63 32.68 15.86 22.48 27.21 69.97 19.66 -0.59%
SnapKV0.9 70.20 27.90 20.08 25.38 52.72 24.04 43.93 33.92 15.56 22.49 27.65 70.18 19.05 -1.12%
H2O0.5 31.08 2.90 16.47 17.99 42.37 23.14 43.05 32.79 15.77 22.79 26.24 69.83 19.18 -17.63%
SLM0.5 3.41 6.03 8.84 18.75 44.68 20.94 37.73 31.20 15.29 21.80 25.93 67.79 19.12 -24.73%
SnapKV0.5 12.96 7.81 16.87 19.52 42.84 23.40 43.93 33.98 15.94 22.67 26.07 69.66 19.19 -17.03%
Mistral-7B-Instruct Full 33.36 29.24 6.83 20.82 39.68 28.74 37.80 33.87 22.88 22.19 22.94 86.87 16.08 100%
Oursk=1 31.54 29.02 6.71 20.81 39.80 27.20 38.93 33.46 22.15 22.39 22.67 86.81 15.33 -1.50%
Budget 89.6% 90.5% 84.2% 92.3% 84.1% 78.0% 97.9% 86.7% 74.4% 84.3% 89.1% 87.2% 89.4% 86.75%
H2O0.9 19.18 24.11 6.96 20.61 39.74 27.36 38.38 35.23 23.24 22.13 23.41 86.46 16.01 -4.29%
SLM0.9 31.16 27.68 5.89 19.88 39.15 27.26 37.66 35.06 21.94 22.31 21.73 86.63 13.67 -4.44%
SnapKV0.9 27.60 25.67 7.10 20.23 39.76 27.43 38.27 35.66 22.99 22.16 23.31 86.46 16.07 -1.90%
H2O0.5 2.50 8.26 5.89 20.28 38.53 27.62 36.97 34.31 23.17 22.32 22.60 86.52 16.45 -14.31%
SLM0.5 2.43 6.47 1.14 18.27 32.76 24.91 33.80 32.73 20.68 21.48 18.54 86.45 13.58 -27.61%
SnapKV0.5 3.03 8.93 6.83 19.03 39.22 27.15 38.27 34.60 22.94 21.93 22.68 86.31 16.18 -13.41%
Qwen2.5-7B-Instruct Full 88.02 31.70 29.85 24.55 61.43 20.81 43.17 47.15 30.70 23.64 24.24 87.64 2.44 100%
Oursk=1 88.02 31.25 30.39 24.48 60.01 20.66 42.74 47.20 29.56 22.64 24.04 87.65 3.58 +2.63%
Budget 90.7% 88.8% 86.5% 69.2% 84.1% 65.1% 69.2% 73.3% 64.4% 70.6% 85.4% 56.6% 84.4% 76.02%
H2O0.9 83.47 26.56 30.25 24.40 61.29 20.85 43.26 47.90 30.73 23.42 24.35 87.57 2.45 -1.46%
SLM0.9 88.93 30.80 28.25 24.30 60.91 19.86 42.54 46.02 28.75 23.06 23.42 87.03 3.05 -0.41%
SnapKV0.9 80.14 30.13 28.11 24.30 61.26 20.90 43.31 47.81 30.69 23.90 24.29 87.57 2.55 -1.01%
H2O0.5 34.12 14.73 20.88 23.10 59.69 21.59 43.37 47.60 30.81 21.00 22.95 85.54 2.24 -13.47%
SLM0.5 3.26 1.34 10.58 23.75 49.87 19.29 36.45 42.42 25.33 21.24 22.77 87.33 3.40 -23.56%
SnapKV0.5 20.77 12.28 22.76 22.25 60.21 20.93 43.22 47.45 30.61 23.76 22.70 87.57 2.21 -14.39%
Table 1: Performance of DBudgetKV and its comparison with three KV cache pruning models on 13 datasets. Bold numbers indicate the best results aside from full-cache. Italics represent the real budget utilized by DBudgetKV. The correspondence between abbreviations and their full names of tasks and datasets can be found in Appendix. Avg. calculates the mean ratio of the model’s performance using different KV cache compression methods to its performance with the full cache. All average results (except for budget) are adjusted by subtracting 1 to provide a more intuitive understanding of the effectiveness of different pruning methods.
Backbones
Our experiments are conducted with three LLMs: Llama3-Instruct with size of 8B/70B, Mistral-7B-Instruct-V0.3, and Qwen2.5-Instruct with size of 7B/32B/72B. We implement DBudgetKV upon the codebase of SnapKV [^1] https://github.com/FasterDecoding/SnapKV. For the reduced attention matrix $A^{\prime}$, we set $k=1$ in practice (ablation provided in Sec. 4.4), and set $m=4$ for importance ranking.
Datasets
For comprehensive evaluation, we evaluate DBudgetKV using datasets of both short and long context length. For tasks of relatively short inputs, the models are assessed across mathametics, sicence, and commonsense reasoning on multiple datasets including GSM8K Cobbe et al. (2021), GPQA Rein et al. (2023), TheoremQA Chen et al. (2023), TruthfulQA Lin et al. (2022), and CoQA Reddy et al. (2019). For long-context tasks, we adopt tasks and datasets from Longbench Bai et al. (2024). Appendix C provides a detailed description and statistics of these datasets along with how they are utilized in experiments.
Evaluation Protocol
Our primary objective is to evaluate the capability of DBudgetKV to compress KV cache while striking for a full-cache performance. To this end, we compare DBudgetKV with the full KV cache performance and calculate the average compression ratio per dataset. Additionally, we select three previous KV cache compression methods with fixed budget size, including: Heavy Hitter Oracle (H2O) Zhang et al. (2023), StreamingLLM (SLM) Xiao et al. (2024), and SnapKV Li et al. (2024). We compare with two budgets for these three methods at 90% and 50% compression ratio respectively.
Apart from the main experiments, we further evaluate various design factors of DBudgetKV via ablation studies, such as different strategies of token ranking importance, the selection of $k$ for the attention matrix, and the necessity to retain bottom layers. Finally, we also provide efficiency analysis on the time and space overhead comparing against other KV cache pruning methods in Sec. 4.5.
4.2 Main Results
The main results are shown in Table 1 according to our evaluation protocol, where DBudgetKV demonstrates strong advantages with robust performance across diverse tasks and models:
• DBudgetKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types. Interestingly, with Llama3-8B and Qwen2.5-7B, DBudgetKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache. With Mistral, DBudgetKV also manages to achieve nearly 20% KV cache compression with only a 1.5% performance reduction. These results indicate that our proposed method can be practically deployed in real-world scenarios without worrying domain-specific budget thresholds. In stark contrast, previous methods achieve a consistent full-cache performance only when manually determined a high budget ratio, e.g. 90%. However, when the budget is reduced, e.g. 50%, the degradation can become severe on certain datasets, distinct from DBudgetKV that automatically adjusts the pruning to always attain full performance.
Methods GSM8K CoQA NQA Musique QMSum
Llama3-70B-Instruct
Full 89.69 60.36 27.15 29.31 22.52
Oursk=1 89.76 60.38 26.94 28.88 22.30
Budget 93.3% 77.2% 58.7% 60.8% 52.9%
Qwen2.5-32B-Instruct
Full 91.36 58.03 24.78 40.04 22.84
Ours 91.81 57.29 22.65 40.54 22.52
Budget 91.6% 85.0% 68.2% 74.4% 76.8%
Qwen2.5-72B-Instruct
Full 90.22 54.14 24.36 42.13 23.93
Ours 90.30 54.19 24.10 41.70 23.31
Budget 90.9% 87.6% 63.4% 65.0% 68.6%
Table 2: Performance of LLMs with scales of 70B, 34B, and 72B across five datasets using DBudgetKV.
Methods GSM8K CoQA NQA Musique QMSum
Full 75.28 52.74 24.06 14.77 22.27
Post=1% Performance Using Different $k$
$k=1$ 76.50 52.86 23.44 15.96 21.95
Budget 93.2% 81.5% 48.7% 43.7% 15.0%
$k=1\%n$ 76.19 52.87 23.17 15.40 21.94
Budget 95.1% 94.8% 77.3% 77.2% 59.5%
$k=5\%n$ 75.59 52.75 21.62 13.71 21.43
Budget 98.6% 96.2% 45.1% 30.0% 33.1%
$k=10\%n$ 76.72 52.85 9.74 13.67 21.11
Budget 96.3% 96.8% 32.0% 19.7% 29.3%
$k=1$ Performance with Different Ranking Method
Attnt=1% 2.35 43.68 13.37 9.58 17.62
Budget 20.0% 9.7% 6.4% 6.4% 6.4%
Attnt=0.01% 54.66 51.58 17.45 14.79 20.90
Budget 61.7% 28.6% 6.7% 7.9% 7.1%
Posk=1 Performance with Different Threshold $t$
${t=0.1\%}$ 76.35 52.84 22.73 17.04 21.90
Budget 99.2% 97.0% 77.2% 85.0% 73.9%
${t=10\%}$ 26.23 49.69 18.35 15.16 18.81
Budget 55.8% 37.2% 15.27% 18.5% 6.5%
Table 3: Performance comparison with different $k$ values, ranking methods and universal thresholds.
• DBudgetKV natually reflects the difficulty of the generation task. As in Table 1, the dynamic budget ratio is high on Math&Science datasets (over 90%), while much lower on QA or Summarization datasets (as low as 15%). This observation is in line with our intuition, where inference on concise but hard tasks, such as math problems, requires more context and more precise calculation, resulting in higher budget allocation.
• Besides the dynamic compression, DBudgetKV also outperforms the three 90%-budget baselines, while itself uses less than 90% budget, and achieves the highest score in 20 out of 39 comparisons. This suggests that by taking into account both the position bias and attention scores, our method accompolishes more effective pruning than the baselines that only consider one aspect, highlighting the importance of integrating both dimensions for effective KV cache compression.
Methods GSM8K CoQA NarrativeQA Musique QMSum TriviaQA Avg.
Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune
Llama3-8B-Instruct
Full 4.693 0.289 3.441 3.659 5.458 2.717 3.376
Oursk=1 4.638 0.034 0.331 0.033 3.159 0.255 3.658 0.264 5.330 0.241 2.684 0.211 3.300 0.173
H2O0.5 4.686 0.020 0.290 0.018 3.243 0.264 3.679 0.266 5.398 0.249 2.612 0.237 3.318 0.176
SLM0.5 6.071 0.006 0.301 0.030 3.230 0.257 3.784 0.259 5.454 0.234 2.628 0.208 3.578 0.166
SnapKV0.5 5.874 0.021 0.285 0.019 3.307 0.265 3.721 0.266 5.519 0.264 2.650 0.265 3.559 0.183
Llama3-70B-Instruct
Full 17.504 1.167 6.285 6.565 13.993 5.302 8.469
Oursk=1 15.975 0.154 1.253 0.290 6.042 2.345 7.412 2.352 13.835 3.028 5.224 1.602 8.290 1.623
H2O0.5 19.062 0.114 1.059 0.240 6.788 2.298 7.082 2.307 14.360 3.712 5.566 1.308 8.986 1.663
SLM0.5 22.079 0.150 1.138 0.229 6.804 2.069 7.107 2.053 14.734 3.700 5.576 1.367 9.573 1.595
SnapKV0.5 16.517 0.117 1.116 0.241 6.795 2.474 7.085 2.468 13.846 2.793 5.562 1.083 8.487 1.529
Table 4: The average inference time and pruning time of Llama3 with size of 8B/70B across six datasets, measured in seconds, with lower values indicating better performance.
4.3 DBudgetKV’s Generalization Ability
We next conduct experiments to examine whether our design and hyperparameters can be generalized to LLMs of larger sizes. On five datasets in Table 2, DBudgetKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same DBudgetKV setting. Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% lossless compression at its peak, validating the robustness and versatility for its general applicability.
4.4 Ablation Study
The ablation study is conducted to explore the impact of various configurations of DBudgetKV. We select Llama-3-8B-Instruct and perform experiments across five datasets presented in Table 3.
Attention Matrix Reduction
The reduced attention matrix $A^{\prime}$ in Sec. 3.2 aggregates attention scores from the last $k$ positions. The upper part of Table 3 illustrates the model’s performance and the actual budget when setting $k=1$, $1\%n$, $5\%n$, and $10\%n$ (n being the number of input tokens). It is evident that setting $k$ as $1$ achieves a significantly reduced budget, thus a higher compression ratio, with almost no change in performance compared to $1\%n$ and $5\%n$. On the other hand, while $10\%n$ can compress more KV cache, it fails to maintain performance (for instance, on NarrativeQA, the former achieves a performance of 9.74 using 32% of the budget, whereas the latter scores 21.44 using 48.7% of the budget). What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last token, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length. The advantages of both high efficacy and low overhead make $k=1$ a solid design choice for calculating the norm metric.
Importance Ranking Strategies
DBudgetKV simply regards positions as the significance of tokens, while another intuitive strategy is to utilize attention scores to rank their importance. We investigate its potentials through ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O. The results are shown in the middle of Table 3: under the same experimental settings, the attention-based ranking struggles to maintain the model’s performance. Though it achieves a high compression ratio, the performance is not stable with possible severe degradation, which deviates from our objective. Therefore, the position-based strategy is more suited for our goal.
Attention Norm Threshold
As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in the bottom part of Table 3. Intuitively, a larger threshold allows for a higher compression ratio, while a smaller threshold does the opposite. The conclusion is clear that when the threshold is set to a smaller 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement. Conversely, when the threshold is set to 10%, it prunes more KV cache, but the model’s performance significantly deteriorates. Thus, we deem 1% as a reasonable threshold for universal uses across models and tasks.
Appendix D presents additional results and analysis from the ablation study.
4.5 Efficiency Analysis
In this section, a quantitative study is conducted on whether DBudgetKV improves inference time, apart from the space reduction from KV cache pruning. We compare the time usage on six datasets with Llama3-8B/70B using DBudgetKV, along with three baselines with the 50% budget setting. Table 4 reports the average generation time after the prefilling stage of each sample (Overall), as well as the time needed to complete pruning (Prune) before the generation.
From the results, DBudgetKV achieves the best performance in 8 out of 12 comparisons of overall generation time, and achieves the best average time for both 8B and 70B LLMs, suggesting that although DBudgetKV requires a slightly longer time for pruning itself compared to three other approaches, the total generation speed of DBudgetKV is evidently advantageous. It also performs consistently across models of different scales, underscoring its time efficiency.
5 Conclusion
In this study, we introduce an innovative KV cache compression objective designed to approximate the full-cache performance, independent of specific inputs, while optimizing resource utilization through targeted KV cache pruning. Our approach, termed DBudgetKV, employs a straightforward yet effective two-step process for each layer of LLMs. Tokens are initially ranked solely based on their positions, followed by the eviction of the least significant tokens as determined by the norm value of the reduced attention matrix. Comprehensive experiments conducted across diverse datasets, encompassing a variety of tasks and context lengths, demonstrate that DBudgetKV achieves nearly lossless compression, while with notable space and time reduction, successfully fulfilling our objective for greater practical values.
Limitations
The main limitation of DBudgetKV stems from the gap between its current compression budget and the true optimal budget. This can be seen in Table 1 that for certain scenarios, e.g. QMSum with Mistral-7B, DBudgetKV reaches 84.3% budget while 50% budget is also viable with no performance degradation. We regard this gap as room for improvement in future work, facilitating more aggressive KV cache compression while ensuring the full-cache performance.
Another limitation of DBudgetKV is that though it demonstrates almost lossless compression, we rely on empirical experiments and there is no hard guarantee on the full-performance constraint. In Table 1, while both Llama3-8B and Qwen2.5-7B even surpass the full-cache performance, DBudgetKV with Mistral-7B shows trivial degradation of 1.5%. More robust methods could be developed as future work to further strengthen this constraint.
Appendix A Full Algorithm of DBudgetKV
Algorithm 1 DBudgetKV
  Input: Prompt, Threshold $t$
  Output: Compressed KV Cache
  Create Empty List $K_{c},V_{c}$
  for Transformer Layer $L_{i}$ in LLM do
     $Q^{i},K^{i},V^{i}\leftarrow L_{i}(\text{Prompt})$
     $R^{i}\leftarrow$ Postion-Based Importance Rank
     $A_{last}^{i}\leftarrow\operatorname{Attention}(Q^{i}[\dots,-1,:],K^{iT})$
     $F_{b}^{i}\leftarrow\operatorname{Frobenius}(A_{last}^{i})$
     $A_{last}^{i}\leftarrow\operatorname{Square}(A_{last}^{i})$
     Reorder $A_{last}^{i}$ by Rank $R$
     $A_{cumsum}^{i}\leftarrow\operatorname{Cumsum}(A_{last}^{i})$
     $A_{cumsum}^{i}\leftarrow\operatorname{Sqrt}(A_{cumsum}^{i})$
     $A_{ratio}^{i}\leftarrow(F_{b}^{i}-A_{cumsum}^{i})/F_{b}^{i}$
     $\text{Index}^{i}\leftarrow\operatorname{Max}(\operatorname{Where}(A_{ratio}^{i}<=t))$
     $K_{c}^{i}\leftarrow\operatorname{Compress}K^{i}\text{by}R^{i}[I:]$
     $V_{c}^{i}\leftarrow\operatorname{Compress}V^{i}\text{by}R^{i}[I:]$
     Append $K_{c}^{i},V_{c}^{i}$ to $K_{c},V_{c}$
  end for
  return  $K_{c},V_{c}$
Appendix B Freeze the first two layers of LLM
In this section, we elaborate the necessity to freeze the first and second layers in DBudgetKV as mentioned in Section 3. In our preliminary studies, we apply DBudgetKV to Llama2-7B-Chat and conduct experiments on GSM8K and CoQA. We first perform a case study, followed by a comparison with the effects of not freezing the bottom layers.
As shown in Table 5, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence. However, the budget exceeds 90% in both cases. By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 0 and 1 are relatively low, while the budgets from layer 2 to layer 30 become very high, with the last layer being low again. We hypothesize that the model is still encoding the global semantics in its early layers, thus not yet able to identify truly important tokens in the first two layers, of which the attention distribution is relatively uniform. This phenomenon has also been observed by early works on Transformers analysis Ethayarajh (2019); Garí Soler and Apidianaki (2021). Not freezing early layers may lead to early eviction of important tokens, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
Based on the above case study, we attempt to freeze certain layers of the model and explore the optimal freezing configuration. We continue to validate the results of freezing different layers, and the resuls are shown in Table 6. We can observe that freezing the first two layers achieves a balance between model performance and budget. Moreover, freezing the last layer does not significantly enhance the model’s performance and instead leads to an increase in budget. DBudgetKV ultimately opts to begin KV cache compression from the 2nd layer of the model.
GSM8K
Input
A robe takes 2 bolts of blue fiber a-
nd half that much white fiber. How
many bolts in total does it take?
Budget
Layer 0: 67.71
Layer 1: 82.29
Layer 2$\sim$30: 95.83
Layer 31: 37.50
Avg.: 93.90
Output
I have determined by answering the
answer to the format of bolts bolts b-
olts…(repeat)
Ground-Truth 3
CoQA
Input
You are given a story and a question.
Answer the question as concisely as
you can…Question: What color was
Cotton?
Budget
Layer 0: 42.14
Layer 1: 42.54
Layer 2$\sim$30: 99.19
Layer 31: 79.64
Avg.: 95.03
Output
Question: Question: What is the que-
stion: What is the question:
Groud-Truth White
Table 5: Case Study.
Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
Table 6: Performance of Llama2-7B-Chat on GSM8K and CoQA using DBudgetKV with different frozen layers.
Appendix C Datasets Used in Experiments
In this section, we provide a comprehensive overview of all the tasks and datasets utilized in the experiments in this paper.
Math & Science
This task evaluates the model’s ability to tackle mathematical and scientific problems. By directly inputting questions and comparing the model’s output with the correct answers, we calculate the model’s Accuracy on these datasets: GSM8K is a dataset for evaluating model’s math-solving skills, featuring 8,000 elementary-level math word problems requiring basic arithmetic and reasoning. GPQA tests model’s understanding of physics concepts and problem-solving across various topics, assessing scientific reasoning abilities. TheoremQA evaluates model’s grasp and application of mathematical theorems, ranging from simple applications to complex proofs, testing advanced math skills.
Commonsense Reasoning (CR)
This task evaluates model’s ability to make deductions and understand everyday situations using implicit knowledge and logical inference. TruthfulQA (ThQA) evaluates model’s ability to generate accurate and truthful responses, testing models on distinguishing fact from fiction, especially in areas prone to misconceptions. We use BLEU as the metric. CoQA assesses model’s ability to understand and respond to questions in a conversational context, focusing on maintaining coherence and context throughout a dialogue. We use F1 Score as the metric.
Single Document QA (Single-Doc QA)
This task assesses the model’s reading comprehension skills when dealing with a single, extended document. NarrativeQA Kociský et al. (2018) is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes. Qasper Dasigi et al. (2021) is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information. We employ F1 Score as the metric for above two datasets.
Multi-Document QA (Multi-Doc QA)
This task evaluates the model’s reading comprehension capabilities across multiple extended documents. 2WikiMultiHopQA (2WKMQA) Ho et al. (2020) is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles. MuSiQue Trivedi et al. (2022) evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately. We leverage F1 Score as the metric for above two datasets.
Methods GSM8K CoQA NQA Musique QMSum
Full 75.28 52.74 24.06 14.77 22.27
Post=1% Performance Using Different $k$
$k=1$ 76.50 52.86 23.44 15.96 21.95
Budget 93.2% 81.5% 48.7% 43.7% 15.0%
$k=1\%n$ 76.19 52.87 23.17 15.40 21.94
Budget 95.1% 94.8% 77.3% 77.2% 59.5%
$k=2\%n$ 76.19 52.82 22.87 14.34 21.74
Budget 96.0% 96.1% 69.8% 62.0% 47.1%
$k=3\%n$ 75.82 52.80 23.03 13.61 21.56
Budget 98.6% 95.7% 60.3% 46.0% 39.5%
$k=4\%n$ 75.21 52.76 22.96 13.45 21.47
Budget 98.8% 95.8% 51.9% 35.2% 35.4%
$k=5\%n$ 75.59 52.75 21.62 13.71 21.43
Budget 98.6% 96.2% 45.1% 30.0% 33.1%
$k=6\%n$ 75.66 52.81 21.95 14.33 20.94
Budget 98.5% 96.6% 40.4% 25.8% 31.6%
$k=7\%n$ 76.57 52.88 21.18 13.66 20.94
Budget 98.0% 96.7% 37.3% 22.5% 30.6%
$k=8\%n$ 76.35 52.86 20.36 13.75 21.11
Budget 97.4% 96.7% 35.1% 20.9% 30.1%
$k=9\%n$ 76.65 52.84 19.99 13.66 20.76
Budget 96.8% 96.8% 33.4% 19.9% 29.5%
$k=10\%n$ 76.72 52.85 9.74 13.67 21.11
Budget 96.3% 96.8% 32.0% 19.7% 29.3%
Table 7: Performance comparison with different $k$values, ranking methods and universal thresholds.
Summarization
This task examines the model’s ability to comprehend and summarize lengthy documents. QMSum Zhong et al. (2021) is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions. Multi-News (M-News) Fabbri et al. (2019) is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic. We use Rouge-L as the metric for above two datasets.
Few-Shot Learning (FSL)
This task assesses the model’s few-shot learning capabilities. TriviaQA Joshi et al. (2017) is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall. We use F1 Score as the metric.
Code
This task evaluates the model’s ability to complete and generate code. LCC Guo et al. (2023) is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures. We use Edit Sim as the metric.
Appendix D Ablation
In this section, we present additional ablation study results. By setting various values for $k$, we expand upon the results shown in Table 3. The expanded results are displayed in Table 7. These experiments facilitate a deeper understanding of how different parameter settings impact model performance and provide a basis for optimizing parameter selection.
As shown in Table 7, setting $k=1$ not only conserves pruning time but also achieves better model performance with a reduced budget.

### v2

@@
 To alleviate memory burden during inference of large language models (LLMs), numerous studies have focused on compressing the KV cache by exploring aspects such as attention sparsity.
-However, these techniques often require a pre-defined cache budget; as the optimal budget varies with different input lengths and task types, it limits their practical deployment accepting open-domain instructions.
+These techniques are often designed with a pre-defined KV budget; however, as the optimal budget varies by different input lengths and task types, the existence of a fixed budget could result in inconsistent performance accepting inputs of diverse domains.
 To address this limitation, we propose a new KV cache compression objective: to always ensure the full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible.
@@
 Furthermore, our method is easy to integrate within LLM inference, not only optimizing memory space, but also showing reduced inference time compared to existing methods.
-DBudgetKV: Dynamic Budget in KV Cache Compression
-for Ensuring Optimal Performance
-Xuanfan Nithanks: Equal contributions.
-Partial work was done during Xuanfan’s internship at WeChat AI.
-Liyan Xu11footnotemark: 133footnotemark: 3   Chenyang Lyu  Longyue Wang Mo Yu33footnotemark: 3   Lemao Liu33footnotemark: 3   Fandong Meng33footnotemark: 3   Jie Zhou33footnotemark: 3   Piji Lithanks: Corresponding author.
-33footnotemark: 3Pattern Recognition Center, WeChat AI xuanfanni@gmail.comliyanlxu@tencent.comlipiji.pz@gmail.com
 1 Introduction
-As most large language models (LLMs) follows autoregressive generation Minaee et al.
-(2024); Chang et al.
-(2023); OpenAI (2023), they rely on KV Cache to store intermediate states during inference.
-Specifically, LLMs access the cached key and value vectors of past tokens during the attention calculation, thereby speeding up inference by re-using these KV cache Vaswani et al.
-(2017); Ainslie et al.
-(2023); Shazeer (2019).
-However, as the model size and the input length increase, the memory required to maintain the KV cache also grows proportionally.
+Most large language models (LLMs) follows autoregressive generation, where they rely on KV Cache to store intermediate states during inference [1, 2, 3].
+Specifically, LLMs access the cached key and value vectors of past tokens during the attention calculation, thereby speeding up inference by re-using these KV cache [4, 5, 6].
+However, when the model size and the input length increase, the memory required to maintain the KV cache also grows proportionally.
 Since these vectors are usually stored within GPU memory, managing the KV cache efficiently has become crucial to mitigate the overall memory consumption and computational overhead during LLM inference.
-For instance, Llama3 8B model Dubey et al.
-(2024) requires 1 GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
-Consequently, recent works have proposed to effectively reduce KV cache through pruning, leveraging the sparsity of the attention mechanism: certain positions are more pivotal during the generation process, while those less important ones could be pruned with minimal performance degradation; thus, the full KV cache is not always necessary.
-These observations have led to several optimization methods, such as H2O Zhang et al.
-(2023), ScissorHands Liu et al.
-(2023), StreamingLLM Xiao et al.
-(2024), and FastGen Ge et al.
-(2024), where they aim to identify and retain only the most salient token positions, discarding less critical ones based on certain pruning criteria.
-Alongside, other orthogonal paradigms have also been proposed without hard-pruning, such as KVMerger Wang et al.
-(2024) and D2O Wan et al.
-(2024) that merge KV vectors, and MLA DeepSeek-AI et al.
-(2024) that operates attention in the latent space.
-While previous pruning methods have successfully demonstrated that a decent size of KV cache could be discarded in practice, there exists one common limitation that has not been addressed yet: these pruning techniques typically require a pre-defined KV cache budget, of which its optimal threshold could vary significantly according to specific tasks or inputs.
-Hence, in real-world scenarios, one would have to carefully tune many budget thresholds for diverse domains, or set a uniform threshold but likely suffering large degradation on certain tasks.
-We reckon that neither choice is a satisfactory solution, hindering the widespread adoptation of KV pruning techniques.
-To further illustrate, Table 1 shows that when the budget is set to 50%, Llama3-8B-Instruct using H2O achieves over 98% of the full-cache performance on NarrativeQA Kociský et al.
-(2018), whereas under the same conditions on GSM8K Cobbe et al.
-(2021), a widely-used math benchmark, the performance drastically falls to less than 42% of the full-cache performance.
-To address the aforementioned limitation, in this work, we introduce a new objective for KV cache compression: our method aims to strike towards full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible.
+For instance, Llama3 8B model [7] requires 1 GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
+Consequently, recent works have proposed to effectively reduce KV cache, leveraging the sparsity of the attention mechanism: certain positions are more pivotal during the generation process, while those less important ones could be pruned with minimal performance degradation; thus, the full KV cache is not always necessary.
+These observations have led to several optimization methods, such as H2O [8], ScissorHands [9], StreamingLLM [10], and FastGen [11], where they aim to identify and retain only the most salient token positions, discarding less critical ones based on certain pruning criteria.
+Apart from pruning, other paradigms have been proposed as well to reduce KV consumption, such as KVMerger [12] and D2O [13] that merge or compress KV vectors instead of hard-pruning.
+While these studies have successfully demonstrated that a decent portion of KV cache can be discarded in practice, there remains one common issue that has yet to be addressed: these pruning techniques typically require a pre-defined KV cache budget, of which its optimal threshold could vary significantly according to specific tasks or inputs.
+In real-world scenarios, one would need to carefully tune budget thresholds optimized for different domains; otherwise, setting a uniform threshold in diverse domains may potentially bring large degradation on certain tasks.
+We reckon the need for a pre-defined threshold as an overlooked limitation that can lead to unstable and inconsistent performance in practice.
+Table 1: With pre-defined cache budgets, KV pruning methods on Llama3 exhibit inconsistent performance across domains (percentage relative to full-cache scores).
+Methods GSM8K GPQA CoQA
+Budget = 90%
+H2O 99.1% 96.9% 100.2%
+SLM 96.7% 99.2% 100.2%
+SnapKV 93.3% 96.1% 99.9%
+Budget = 20%
+H2O 3.2% 8.5% 96.5%
+SLM 4.3% 11.5% 76.1%
+SnapKV 1.7% 3.9% 98.6%
+To further illustrate, Table 1 shows that when retaining 90% KV cache, Llama3-8B-Instruct could achieve near full-cache performance across three domains.
+However, the performance is shown inconsistent with 20% budget; especially on the math benchmark GSM8K [14], all three KV pruning methods drastically fall to less than 5% relative to the full-cache performance.
+To address the aforementioned limitation, in this work, we propose a new objective for KV cache compression: our method aims to strike towards full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible.
 Towards this objective, we propose an adaptive pruning method, dubbed DBudgetKV, which Dynamically adjusts KV cache retention per input, without the need of pre-defined memory Budget.
-Particularly, DBudgetKV automatically identifies input-dependent KV cache pruning that leads to minimal performance degradation.
+In particular, DBudgetKV automatically identifies input-dependent KV cache pruning that leads to minimal performance degradation.
 A key implication is that the method tends to realize a higher compression ratio for simpler tasks, while allocating more cache resources for complicated tasks.
-Our method features a two-step process: 1) after the prefilling stage, rank all input tokens according to an importance metric; 2) discard the KV cache of each token sequentially, until hitting a stopping criterion.
-In choosing the importance metric and the stopping criterion, DBudgetKV takes into account two important factors from previous works: i) position bias, where tokens at the beginning and the end are generally more significant than those in the middle Xiao et al.
-(2024); Jiang et al.
-(2024b); ii) attention score, where tokens receiving higher scores usually contribute more information to subsequent generation Minaee et al.
-(2024); Chang et al.
-(2023); Zhang et al.
-(2023).
-We then propose DBudgetKV of empirical efficacy and low overhead.
-Concretely, tokens are ranked solely by their positions, with no additional computation involved.
-The ranking strategy signals high significance to the beginning and the end positions, where the initial tokens are the most important, followed by the remaining tokens in a reversing order.
-Next in the pruning process, we design an attetion-based metric to determine the stopping condition, such that the pruning halts when this metric signals that the remaining KV cache is unlikely to match the full-cache capacity, thereby fulfilling our objective to approximate full performance with dynamic cache pruning.
-More specifically, our attention metric compares the difference between the Frobenius form of the full attention matrix and the reduced matrix by setting pruned positions’ score to zero.
-When the difference exceeds a universal threshold, the process is halted.
-Especially, this metric empirically correlates well with the generation performance regardless of inputs, serving as an estimator to bridge the pruning state and the subsequent generation.
-Furthermore, the entire process is implemented efficiently through PyTorch’s built-in operators, where we show that DBudgetKV not only optimizes memory space, but also acheives time reduction compared to the full-cache inference and three KV pruning methods.
+Our method features a two-step process: (1) after the prefilling stage, rank all input tokens according to an importance metric; (2) discard the KV cache of each token sequentially, until hitting a stopping criterion.
+In choosing the importance metric and the stopping criterion, DBudgetKV takes into account two important factors from previous works: i) position bias, where tokens at the beginning and the end are generally more significant than those in the middle [10, 15]; ii) attention score, where tokens receiving higher scores usually contribute more information to subsequent generation [1, 2, 8].
+To eliminate the pre-defined KV budget while always targeting full-cache performance, the stopping criterion becomes crucial, such that it should signal whenever the remaining KV cache is unlikely to match the full-cache capacity, thereby fulfilling the objective to approximate full performance with dynamic cache pruning.
+In DBudgetKV, we design a simple attention-based metric as the stopping condition: the metric compares the difference between the Frobenius form of the full attention matrix and the reduced attention matrix of which pruned positions’ scores are set to zero.
+When the difference exceeds a universal threshold, the pruning process halts.
+In practice, we find that this metric empirically correlates well with the full-cache performance regardless of input types, serving as an estimator to bridge the pruning state and the subsequent generation.
 Figure 1: The overall workflow of DBudgetKV in Section 3.2.
-Initially, tokens are ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix.
-The KV cache for the remaining tokens are then preserved.
-Experimental results on 13 datasets varying diverse context lengths and tasks, e.g.
-mathematical and commonsense reasoning, reading comprehension and coding, demonstrate that DBudgetKV achieves our objective effectively and robustly.
-The resulting inference is on par or even surpasses the full-cache performance with multiple LLMs of different sizes, including Llama3 Dubey et al.
-(2024), Qwen2.5 Yang et al.
-(2024), and Mistral Jiang et al.
-(2024a).
-Meanwhile, the dynamic pruning of DBudgetKV allocates relatively high budget on math benchmarks but attains high compression ratio by up to 85% on comprehension tasks.
-The average budget size reaches 63.7% using Llama3-8B.
-The overall results suggest that DBudgetKV is able to accomplish lossless KV cache pruning with strong generalizability.
-Besides, it does not rely on specific model architectures and can coexist with other KV cache optimization techniques.
-We further conduct more experiments for in-depth analysis and ablation studies, depicting the rationality of our various design choices and hyperparameters.
-Our contributions can be summarized below:
-  • To the best of our knowledge, we are the first to propose the objective that ensures full-cache performance while maximizing the KV cache pruning, which holds greater practical values and deployment potentials.
-  • We introduce DBudgetKV that employs a straightforward two-step process, where its dynamic pruning is designed with minimal degradation to the generation performance.
+After prefilling, tokens are initially ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix.
+The KV cache for the remaining tokens are then preserved to subsequent generation.
+To examine the performance consistency across diverse domains, our experiments adopt 13 datasets varying diverse context lengths and tasks, e.g.
+mathematical and commonsense reasoning, reading comprehension and coding, demonstrating that DBudgetKV achieves our objective effectively and robustly.
+With dynamic KV budgets, the resulting inference is on par or even surpasses the full-cache performance with multiple LLMs of different sizes, including Llama3 [7], Qwen2.5 [16], and Mistral [17].
+The dynamic pruning of DBudgetKV automatically allocates relatively high cache budget on math benchmarks to prevent degradation, while achieving high compression ratio on comprehension tasks by up to 85%.
+With Llama3-8B, DBudgetKV reaches an average budget ratio of 63.7% with minimal degradation across inputs.
+Overall, empirical results suggest that our proposed method is able to accomplish lossless KV cache pruning with strong generalizability.
+Besides, it does not rely on specific model architectures, and is orthogonal to other KV optimization techniques.
+We further conduct experiments for in-depth analysis and ablation studies, where we validate that each step of DBudgetKV is necessary in achieving the desired performance, depicting the rationality of various design choices and hyper-parameters.
+Our contributions can be summarized as follows:
+  • To the best of our knowledge, we are the first to call for the objective that ensures full-cache performance while maximizing the KV cache pruning, addressing the limitation of inconsistent performance with fixed KV budget thresholds.
+  • We introduce DBudgetKV that employs a straightforward two-step process, where the dynamic pruning is halted with minimal degradation to the generation performance.
   • Extensive evaluation with diverse tasks and models indicates that compared to prior works, our method achieves lossless KV cache compression effectively, efficiently and robustly.
@@
 Recent advancements in KV cache compression have emerged to address memory constraints for model inference.
-Scissorhands Liu et al.
-(2023) observes the repetitive nature of attention patterns and the persistence of token importance.
+Scissorhands [9] observes the repetitive nature of attention patterns and the persistence of token importance.
 It then preserves the most important tokens based on attention scores to achieve KV cache compression.
-H2O Zhang et al.
-(2023) introduces dynamic eviction policies that strategically balance retention of recent tokens and historically significant "heavy hitters".
-StreamingLLM Xiao et al.
-(2024) enables infinite-length sequence processing without fine-tuning by stabilizing attention through landmark tokens.
-SnapKV Li et al.
-(2024) enhances compression efficiency through attention score-based selection and clustering of critical KV positions.
-FastGen Ge et al.
-(2024) proposes an adaptive KV cache management system that customizes retention strategies per attention head.
-However, these methods require setting a fixed memory budget, which makes it difficult to consistently maintain full performance across different domains and datasets, limiting their practical application.
-In contrast, our method imposes no such constraints, targeting full-cache performance while allowing for dynamic budget size.
+H2O [8] introduces dynamic eviction policies that strategically balance retention of recent tokens and historically significant "heavy hitters".
+StreamingLLM [10] enables infinite-length sequence processing without fine-tuning by stabilizing attention through landmark tokens.
+SnapKV [18] enhances compression efficiency through attention score-based selection and clustering of critical KV positions.
+FastGen [11] proposes an adaptive KV cache management system that customizes retention strategies per attention head.
+Other paradigms beyond KV pruning have been proposed along the way, such as KV cache merging [19].
+As prior KV pruning methods typically require setting a pre-defined memory budget, they are by design inherently unable to ensure consistent performance across various domains and datasets, which could prevent their widespread deployment in practice.
+Our method targets to lift such constraint, achieving near full-cache performance and dynamic budget size without the need of additional tuning.
+Attention Pattern
+The study of attention head specialization has paved the way for optimizing memory and computation in Transformer models.
+Early research on encoder architectures shows that some attention heads have persistent positional biases, like consistently focusing on adjacent tokens, indicating redundancy in their learned patterns.
+Rho-1 [20] reveals that the importance of individual heads varies significantly across tasks, suggesting that not all attention computations are equally crucial during inference, and reveal the potential for dynamically pruning or approximating redundant computations.
+Recent works [6, 21] have shifted the focus to decoder-only models, where the quadratic memory overhead of retaining intermediate KV vectors poses a significant challenge.
+Focused Transformer [22] extends the context length of LLMs through a contrastive learning-inspired training process.
+Landmark-Attention [23] uses a “landmark” token to represent each block of the input and trains the attention to use it for selecting relevant blocks, enabling retrieval of blocks directly through the attention mechanism instead of by relying on a separate mechanism.
+Our work leverages attention’s structural regularities, particularly the predictable, position-specific behaviors of certain heads, to dynamically reduce the memory footprint of cached KV states.
+By identifying and exploiting the properties with position bias and attention distribution, we enable efficient inference without compromising model performance, aligning with the broader goal of dynamic KV cache pruning for various inputs.
 3 Methodology
@@
 As mentioned in Sec.
-1, most previous works on KV cache pruning require a pre-defined memory budget beforehand, either by fixing the number of cached positions directly Cai et al.
-(2024); Li et al.
-(2024), or retaining positions by a fixed ratio of input length Zhang et al.
-(2023); Wan et al.
-(2024).
+1, most previous works on KV cache pruning require a pre-defined memory budget beforehand, either by fixing the number of cached positions directly [24, 18], or retaining positions by a fixed ratio of input length [8, 13].
 While these approaches are effective in certain circumstances, they may encounter practical challenges when deployed in real-world scenarios.
-The biggest issue is that the optimal budgets for LLM inference in real-world scenarios are evidently infeasible to enumerate, as they could vary across tasks and domains, especially for open-domain instructions.
+It is forthcoming that the optimal budgets for LLM inference in real-world scenarios are evidently infeasible to enumerate, as they could vary across tasks and domains, especially for open-domain instructions.
 For example, mathematical tasks are typically concise and require inference based on all given conditions, thus many positions may logically contribute to the generation, necessitating a relatively higher memory budget.
-Conversely, for article reading comprehension, LLM may only need a small set of critical positions for the answer generation, featuring a smaller budget in many cases.
-As a side effect, inference with previous KV cache pruning is likely to experience performance instability across different inputs, which can be observed by our analysis in Sec.
+Conversely, for article reading comprehension, LLM may only need a small set of critical positions for the answer generation, featuring a smaller budget.
+As a side effect, inference with fixed KV cache pruning may experience performance instability across different inputs, which can be observed by our analysis in Sec.
 4.2.
-Overall, the practical value is thereby diminished.
-In this work, we resort to a dynamic paradigm that eliminates the need for manually setting a memory budget.
+We regard it as a practical limitation that motivates the inception of this work.
+In addressing this limitation, we resort to a dynamic paradigm that eliminates the need for manually setting a memory budget.
 We introduce a new objective for KV cache compression, aiming to automatically reduce the KV cache as much as possible, while fulfilling the condition that the method should always strike towards full-cache performance.
-In this way, one could utilize such pruning method off-the-shelf, without spending time and resources on tuning usage-specific hyperparameters, which we hope would further advance the research development of KV cache compression techniques.
+In this way, one could utilize such pruning method off-the-shelf, without spending time and resources on tuning usage-specific hyper-parameters, which we hope would further advance the research development of KV cache compression techniques.
 3.2 DBudgetKV
-Towards our objective, we design a novel ranking-based pruning method applied after the prefilling stage, consisting of two steps.
+Towards our objective, we design a novel ranking-based pruning method applied after the prefilling stage, comprising two steps.
 First, all tokens are ranked according to an importance metric per Transformers layer.
 Then, the system keeps removing the KV cache sequentially, until hitting a stopping criterion.
-Ultimately, our proposed approach does not need to fix an optimal budget for storing KV cache.
-Rather, the stopping criterion is tied to the norm of attention matrices that works uniformly regardless of the inputs.
+Ultimately, our proposed approach does not need to fix an optimal budget for storing KV cache; rather, it relies on the stopping criterion to determine dynamically, which should work uniformly regardless of the inputs.
 The motivation behind our two-step procedure comes from two phenomena brought by previous works.
-First, there exists position bias such that positions in the beginning and the end are typically important for the subsequent generation, as observed by StreamingLLM Xiao et al.
-(2024).
-Second, attention scores are usually an effective indicator of the token importance, as leveraged by many pruning techniques Liu et al.
-(2023); Zhang et al.
-(2023); Li et al.
-(2024).
+First, there exists position bias such that positions in the beginning and the end are typically important for the subsequent generation, as observed by StreamingLLM [10].
+Second, attention scores are usually an effective indicator of the token importance, as leveraged by many pruning techniques [9, 8, 18].
 For our objective, we propose to combine the two important properties in our approach, described in details as follows.
@@
 By this strategy, the importance in $X$ is ranked from the highest to the lowest as $\{x_{1},x_{2},\dots,x_{m},x_{n},x_{n-1},\dots,x_{m+1}\}$, denoted as $\widehat{X}$, where $m$ is a chosen hyperparameter that works regardless of specific input sequences.
-It should be noted that our choice on importance ranking is empirically supported rather than theoretically based.
-We have experimented other attention-based importance ranking, which are provided in Sec.
-4.4.
-Our position-based strategy is shown both empirically effective and superior on computational overhead.
+Like previous KV compression works, our choice on importance ranking is empirically supported rather than theoretically based.
+We have also experimented other attention-based importance ranking, which are provided in Sec.
+4.5.
+The choice of position-based ranking is empirically shown both empirically effective and superior on computational overhead.
 Stopping Criterion
-With the initially ranked KV vectors, our method then sequantially decides whether to continue or terminate pruning, which needs an evaluation metric serving as the stopping criterion.
-Since we evidently cannot know the precise impact of a KV cache position before the generation, we require a metric that correlates well with the resulting performance change when discarding a position, serving as a bridge to ensure a minimal degradation upon the full KV-cache performance.
+With the initially ranked KV vectors, our method then sequentially decides whether to continue or terminate pruning, which requires an evaluation metric as the stopping criterion.
+Since one evidently cannot know the precise impact of a KV cache position before the generation, we adopt a metric that empirically correlates well with the performance change when discarding a position, serving as a bridge to ensure a minimal degradation upon the full cache performance.
 Inspired by Devoto et al.
-(2024), we utilize the Frobenius norm of the attention matrix that fits the requirement well empirically.
+[25], we utilize the Frobenius norm of the attention matrix that fits the requirement well empirically.
 For a matrix $M\in\mathbb{R}^{m\times n}$, the Frobenius norm of $M$ is the matrix-version of the L2 norm, expressed as $||M||_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}|M_{i,j}|^{2}}$.
+Figure 2: Preliminary experiments on the relationship between the Frobenius norm difference (in percentage) obtained from randomly masking tokens, and the corresponding model performance on GSM8K.
 We conduct preliminary experiments to validate its reliability.
@@
 As illustrated in Figure 2, there is a certain degree of positive correlation between its Frobenius norm and the model’s performance, which are shown generalized across different domains.
-Figure 2: Preliminary experiments on the relationship between the Frobenius norm difference (in percentage) obtained from randomly masking tokens, and the corresponding model performance on GSM8K.
+The simple Frobenius norm works empirically well in subsequent large-scale evaluation in Sec.
+4.
 As the input sequence length $n$ increases, the time and space overhead for norm calculation also grows significantly by $O(n^{2})$.
 Hence, we next aim to reduce the computational scale from the naive norm calculation.
-We observe that the Frobenius norm of the last row of $A$ also correlates with the model’s performance; equivalently saying, the attention distribution of the latest token is good enough to serve as an estimator.
+We observe that the Frobenius norm of the last row of $A$ also correlates with the model’s performance; equivalently speaking, the attention distribution of the latest token is good enough to serve as an estimator.
 Formally, we utilize the last $k$ rows (latest $k$ tokens) of $A$ and reduce to a single attention vector $A^{\prime}\in\mathbb{R}^{1\times n}$.
@@
 We then use the Frobenius norm of $A^{\prime}$, equivalently its L2 norm, as the evaluation metric, which reduces the overall complexity to $O(n)$.
-Finally, for each position in the ranking order, our method compares the difference between the Frobenius norm of the original attention matrix $A$, and the Frobenius norm of the reduced attention vector $A^{\prime}$ of which the corresponding positions are set to 00.
+Finally, for each position in the ranking order, DBudgetKV compares the difference between the Frobenius norm of the original attention matrix $A$, and the Frobenius norm of the reduced attention vector $A^{\prime}$ of which the pruned positions are set to 00.
 We set a universal threshold as 1%, such that when the new norm exceeds 1% difference from the original norm, it hits the stopping criterion.
-The KV vectors of the remaining positions are then all kept, while the preceding positions would have their KV cache discarded.
+The KV vectors of the remaining positions are then all kept, while the preceding positions have their KV cache discarded.
 In this way, our method achieves dynamic KV cache pruning while targeting to maintain the full-cache performance, assisted by the norm metric.
@@
 We hypothesize that this may be due to the relatively uniform distribution of attention scores in these initial layers, leading to the premature discarding of tokens that are crucial for subsequent layers, ultimately causing the entire model’s inference to collapse.
-A more detailed analysis of this issue is provided in Appendix B.
+A more detailed analysis of this issue is provided in Appendix C.
 Based on this finding, we initiate the KV cache compression operations starting from the model’s 2nd layer.
+Regarding ablation studies are further conducted in Sec.4.5.
 4 Experiments
 4.1 Experimental Settings
+Table 2: Performance of DBudgetKV and its comparison with three KV cache pruning models on 13 datasets.
+Bold numbers indicate the best results aside from full-cache.
+Italics represent the real budget utilized by DBudgetKV.
+The correspondence between abbreviations and their full names of tasks and datasets can be found in Appendix.
+Avg.
+calculates the mean ratio of the model’s performance using different KV cache compression methods to its performance with the full cache.
+All average results (except for budget) are adjusted by subtracting 1 to provide a more intuitive understanding of the effectiveness of different pruning methods.
   Methods Math & Science CR Single-Doc QA Multi-Doc QA Summarization FSL Code   Avg.
@@
 SnapKV0.5 12.96 7.81 16.87 19.52 42.84 23.40 43.93 33.98 15.94 22.67 26.07 69.66 19.19 -17.03%
+H2O0.2 2.43 2.46 6.56 20.67 50.91 23.54 42.00 32.50 15.67 22.16 23.68 69.88 19.13 -23.33%
+SLM0.2 3.21 3.35 8.82 18.02 40.11 20.12 36.45 29.91 15.01 21.11 24.68 66.17 18.69 -28.21%
+SnapKV0.2 1.29 1.12 5.89 20.36 52.00 23.10 42.78 32.98 17.01 22.45 24.14 69.63 19.34 -22.89%
 Mistral-7B-Instruct Full 33.36 29.24 6.83 20.82 39.68 28.74 37.80 33.87 22.88 22.19 22.94 86.87 16.08 100%
@@
 SnapKV0.5 3.03 8.93 6.83 19.03 39.22 27.15 38.27 34.60 22.94 21.93 22.68 86.31 16.18 -13.41%
+H2O0.2 1.52 4.46 4.02 18.45 35.55 27.28 33.45 34.36 23.21 21.80 20.93 85.99 14.82 -23.55%
+SLM0.2 1.21 1.79 0.54 18.35 25.45 24.89 28.39 29.53 16.74 20.60 16.75 80.12 14.27 -35.58%
+SnapKV0.2 1.44 0.20 3.35 17.81 37.44 26.76 36.02 33.90 23.05 22.13 20.88 84.27 16.20 -24.28%
 Qwen2.5-7B-Instruct Full 88.02 31.70 29.85 24.55 61.43 20.81 43.17 47.15 30.70 23.64 24.24 87.64 2.44 100%
@@
 SnapKV0.5 20.77 12.28 22.76 22.25 60.21 20.93 43.22 47.45 30.61 23.76 22.70 87.57 2.21 -14.39%
-Table 1: Performance of DBudgetKV and its comparison with three KV cache pruning models on 13 datasets.
-Bold numbers indicate the best results aside from full-cache.
-Italics represent the real budget utilized by DBudgetKV.
-The correspondence between abbreviations and their full names of tasks and datasets can be found in Appendix.
-Avg.
-calculates the mean ratio of the model’s performance using different KV cache compression methods to its performance with the full cache.
-All average results (except for budget) are adjusted by subtracting 1 to provide a more intuitive understanding of the effectiveness of different pruning methods.
+H2O0.2 5.91 3.57 17.40 21.70 55.39 21.41 40.30 45.84 29.72 23.04 20.84 87.99 2.15 -13.89%
+SLM0.2 1.06 4.24 4.82 23.85 40.92 18.17 29.22 39.27 21.71 20.32 18.94 77.80 1.46 -31.66%
+SnapKV0.2 2.58 7.37 15.13 21.52 57.06 21.40 40.93 46.96 29.87 23.58 20.69 87.10 2.44 -13.26%
 Backbones
@@
 For the reduced attention matrix $A^{\prime}$, we set $k=1$ in practice (ablation provided in Sec.
-4.4), and set $m=4$ for importance ranking.
+4.5), and set $m=4$ for importance ranking.
 Datasets
 For comprehensive evaluation, we evaluate DBudgetKV using datasets of both short and long context length.
-For tasks of relatively short inputs, the models are assessed across mathametics, sicence, and commonsense reasoning on multiple datasets including GSM8K Cobbe et al.
-(2021), GPQA Rein et al.
-(2023), TheoremQA Chen et al.
-(2023), TruthfulQA Lin et al.
-(2022), and CoQA Reddy et al.
-(2019).
-For long-context tasks, we adopt tasks and datasets from Longbench Bai et al.
-(2024).
-Appendix C provides a detailed description and statistics of these datasets along with how they are utilized in experiments.
+For tasks of relatively short inputs, the models are assessed across mathematics, science, and commonsense reasoning on multiple datasets including GSM8K [14], GPQA [26], TheoremQA [27], TruthfulQA [28], and CoQA [29].
+For long-context tasks, we adopt tasks and datasets from Longbench [30].
+Appendix A provides a detailed description and statistics of these datasets along with how they are utilized in experiments.
 Evaluation Protocol
@@
 To this end, we compare DBudgetKV with the full KV cache performance and calculate the average compression ratio per dataset.
-Additionally, we select three previous KV cache compression methods with fixed budget size, including: Heavy Hitter Oracle (H2O) Zhang et al.
-(2023), StreamingLLM (SLM) Xiao et al.
-(2024), and SnapKV Li et al.
-(2024).
-We compare with two budgets for these three methods at 90% and 50% compression ratio respectively.
+Additionally, we select three previous KV cache compression methods with fixed budget size, including: Heavy Hitter Oracle (H2O) [8], StreamingLLM (SLM) [10], and SnapKV [18].
+We compare with three budgets for these three methods at 90%, 50% and 20% compression ratio respectively.
 Apart from the main experiments, we further evaluate various design factors of DBudgetKV via ablation studies, such as different strategies of token ranking importance, the selection of $k$ for the attention matrix, and the necessity to retain bottom layers.
 Finally, we also provide efficiency analysis on the time and space overhead comparing against other KV cache pruning methods in Sec.
-4.5.
+4.4.
 4.2 Main Results
 The main results are shown in Table 2 according to our evaluation protocol, where DBudgetKV demonstrates strong advantages with robust performance across diverse tasks and models:
-• DBudgetKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types.
-Interestingly, with Llama3-8B and Qwen2.5-7B, DBudgetKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache.
-With Mistral, DBudgetKV also manages to achieve nearly 20% KV cache compression with only a 1.5% performance reduction.
-These results indicate that our proposed method can be practically deployed in real-world scenarios without worrying domain-specific budget thresholds.
-In stark contrast, previous methods achieve a consistent full-cache performance only when manually determined a high budget ratio, e.g.
-90%.
-However, when the budget is reduced, e.g.
-50%, the degradation can become severe on certain datasets, distinct from DBudgetKV that automatically adjusts the pruning to always attain full performance.
-Methods GSM8K CoQA NQA Musique QMSum
+Table 3: Performance of LLMs with scales of 70B, 32B, and 72B across five datasets using DBudgetKV.
+GSM8K CoQA NQA Musique QMSum
 Llama3-70B-Instruct
@@
 Budget 90.9% 87.6% 63.4% 65.0% 68.6%
-Table 2: Performance of LLMs with scales of 70B, 34B, and 72B across five datasets using DBudgetKV.
+• DBudgetKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types.
+Interestingly, with Llama3-8B and Qwen2.5-7B, DBudgetKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache.
+With Mistral, DBudgetKV also manages to achieve nearly 20% KV cache compression with only a 1.5% performance reduction.
+These results indicate that our proposed method can be suited for real-world scenarios with no bother by domain-specific budget thresholds.
+In stark contrast, previous methods achieve a consistent full-cache performance only when manually determined a high budget ratio, e.g.
+90%.
+However, when the budget is reduced, e.g.
+50% and 20%, the degradation can become severe on certain datasets, distinct from DBudgetKV that automatically adjusts the pruning to always prioritize full performance.
+• DBudgetKV naturally reflects the difficulty of the generation task.
+As in Table 2, the dynamic budget ratio is high on Math&Science datasets (over 90%), while much lower on QA or Summarization datasets (as low as 15%).
+This observation is in line with our intuition, where inference on concise but hard tasks, such as math problems, requires more context and more precise calculation, resulting in higher budget allocation.
+Thus, our method design serves beyond for inference efficiency, but could also be leveraged for input analysis in a broader scope.
+• Besides the dynamic compression, DBudgetKV also outperforms the three 90%-budget baselines, while itself uses less than 90% budget, and achieves the highest score in 20 out of 39 comparisons.
+This suggests that by taking into account both the position bias and attention scores, our method accompolishes more effective pruning than the baselines that only consider one aspect, highlighting the importance of integrating both dimensions for effective KV cache compression.
+It is worth noting that the scope of this work is not to propose yet another KV cache pruning method.
+Instead, we identify the common limitation of previous KV pruning methods, and call for optimal KV pruning without a pre-defined budget.
+Consequently, the consistent performance across domains after KV pruning is the ultimate goal of our method, instead of achieving the highest compression ratio on certain datasets.
+Under our objective, we leave the further improvement on compression ratio as a promising future research direction.
+4.3 DBudgetKV’s Generalization Ability
+We next conduct experiments to examine whether our design and hyper-parameters can be generalized to LLMs of larger sizes.
+On five datasets in Table 3, DBudgetKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same DBudgetKV setting.
+Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% compression.
+But most importantly, they are able to maintain lossless performance across inputs, validating the robustness and versatility for its general applicability.
+4.4 Efficiency Analysis
+Table 4: The average inference time and pruning time of Llama3 with size of 8B/70B across six datasets, measured in seconds, with lower values indicating better performance.
+GSM8K CoQA NarrativeQA Musique QMSum TriviaQA Avg.
+Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune
+Llama3-8B-Instruct
+Full 4.693 0.289 3.441 3.659 5.458 2.717 3.376
+Oursk=1 4.638 0.034 0.331 0.033 3.159 0.255 3.658 0.264 5.330 0.241 2.684 0.211 3.300 0.173
+H2O0.5 4.686 0.020 0.290 0.018 3.243 0.264 3.679 0.266 5.398 0.249 2.612 0.237 3.318 0.176
+SLM0.5 6.071 0.006 0.301 0.030 3.230 0.257 3.784 0.259 5.454 0.234 2.628 0.208 3.578 0.166
+SnapKV0.5 5.874 0.021 0.285 0.019 3.307 0.265 3.721 0.266 5.519 0.264 2.650 0.265 3.559 0.183
+Llama3-70B-Instruct
+Full 17.504 1.167 6.285 6.565 13.993 5.302 8.469
+Oursk=1 15.975 0.154 1.253 0.290 6.042 2.345 7.412 2.352 13.835 3.028 5.224 1.602 8.290 1.623
+H2O0.5 19.062 0.114 1.059 0.240 6.788 2.298 7.082 2.307 14.360 3.712 5.566 1.308 8.986 1.663
+SLM0.5 22.079 0.150 1.138 0.229 6.804 2.069 7.107 2.053 14.734 3.700 5.576 1.367 9.573 1.595
+SnapKV0.5 16.517 0.117 1.116 0.241 6.795 2.474 7.085 2.468 13.846 2.793 5.562 1.083 8.487 1.529
+Table 5: Performance comparison for ablation studies in Section 4.5, with different $k$ values for attention matrix reduction, ranking methods and universal thresholds.
 Methods GSM8K CoQA NQA Musique QMSum
@@
 Budget 55.8% 37.2% 15.27% 18.5% 6.5%
-Table 3: Performance comparison with different $k$ values, ranking methods and universal thresholds.
-• DBudgetKV natually reflects the difficulty of the generation task.
-As in Table 1, the dynamic budget ratio is high on Math&Science datasets (over 90%), while much lower on QA or Summarization datasets (as low as 15%).
-This observation is in line with our intuition, where inference on concise but hard tasks, such as math problems, requires more context and more precise calculation, resulting in higher budget allocation.
-• Besides the dynamic compression, DBudgetKV also outperforms the three 90%-budget baselines, while itself uses less than 90% budget, and achieves the highest score in 20 out of 39 comparisons.
-This suggests that by taking into account both the position bias and attention scores, our method accompolishes more effective pruning than the baselines that only consider one aspect, highlighting the importance of integrating both dimensions for effective KV cache compression.
-Methods GSM8K CoQA NarrativeQA Musique QMSum TriviaQA Avg.
-Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune
-Llama3-8B-Instruct
-Full 4.693 0.289 3.441 3.659 5.458 2.717 3.376
-Oursk=1 4.638 0.034 0.331 0.033 3.159 0.255 3.658 0.264 5.330 0.241 2.684 0.211 3.300 0.173
-H2O0.5 4.686 0.020 0.290 0.018 3.243 0.264 3.679 0.266 5.398 0.249 2.612 0.237 3.318 0.176
-SLM0.5 6.071 0.006 0.301 0.030 3.230 0.257 3.784 0.259 5.454 0.234 2.628 0.208 3.578 0.166
-SnapKV0.5 5.874 0.021 0.285 0.019 3.307 0.265 3.721 0.266 5.519 0.264 2.650 0.265 3.559 0.183
-Llama3-70B-Instruct
-Full 17.504 1.167 6.285 6.565 13.993 5.302 8.469
-Oursk=1 15.975 0.154 1.253 0.290 6.042 2.345 7.412 2.352 13.835 3.028 5.224 1.602 8.290 1.623
-H2O0.5 19.062 0.114 1.059 0.240 6.788 2.298 7.082 2.307 14.360 3.712 5.566 1.308 8.986 1.663
-SLM0.5 22.079 0.150 1.138 0.229 6.804 2.069 7.107 2.053 14.734 3.700 5.576 1.367 9.573 1.595
-SnapKV0.5 16.517 0.117 1.116 0.241 6.795 2.474 7.085 2.468 13.846 2.793 5.562 1.083 8.487 1.529
-Table 4: The average inference time and pruning time of Llama3 with size of 8B/70B across six datasets, measured in seconds, with lower values indicating better performance.
-4.3 DBudgetKV’s Generalization Ability
-We next conduct experiments to examine whether our design and hyperparameters can be generalized to LLMs of larger sizes.
-On five datasets in Table 2, DBudgetKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same DBudgetKV setting.
-Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% lossless compression at its peak, validating the robustness and versatility for its general applicability.
-4.4 Ablation Study
-The ablation study is conducted to explore the impact of various configurations of DBudgetKV. We select Llama-3-8B-Instruct and perform experiments across five datasets presented in Table 3.
-Attention Matrix Reduction
-The reduced attention matrix $A^{\prime}$ in Sec.
-3.2 aggregates attention scores from the last $k$ positions.
-The upper part of Table 3 illustrates the model’s performance and the actual budget when setting $k=1$, $1\%n$, $5\%n$, and $10\%n$ (n being the number of input tokens).
-It is evident that setting $k$ as $1$ achieves a significantly reduced budget, thus a higher compression ratio, with almost no change in performance compared to $1\%n$ and $5\%n$.
-On the other hand, while $10\%n$ can compress more KV cache, it fails to maintain performance (for instance, on NarrativeQA, the former achieves a performance of 9.74 using 32% of the budget, whereas the latter scores 21.44 using 48.7% of the budget).
-What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last token, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length.
-The advantages of both high efficacy and low overhead make $k=1$ a solid design choice for calculating the norm metric.
-Importance Ranking Strategies
-DBudgetKV simply regards positions as the significance of tokens, while another intuitive strategy is to utilize attention scores to rank their importance.
-We investigate its potentials through ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O.
-The results are shown in the middle of Table 3: under the same experimental settings, the attention-based ranking struggles to maintain the model’s performance.
-Though it achieves a high compression ratio, the performance is not stable with possible severe degradation, which deviates from our objective.
-Therefore, the position-based strategy is more suited for our goal.
-Attention Norm Threshold
-As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in the bottom part of Table 3.
-Intuitively, a larger threshold allows for a higher compression ratio, while a smaller threshold does the opposite.
-The conclusion is clear that when the threshold is set to a smaller 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement.
-Conversely, when the threshold is set to 10%, it prunes more KV cache, but the model’s performance significantly deteriorates.
-Thus, we deem 1% as a reasonable threshold for universal uses across models and tasks.
-Appendix D presents additional results and analysis from the ablation study.
-4.5 Efficiency Analysis
 In this section, a quantitative study is conducted on whether DBudgetKV improves inference time, apart from the space reduction from KV cache pruning.
@@
 It also performs consistently across models of different scales, underscoring its time efficiency.
+4.5 Ablation Study
+The ablation study is conducted to explore the impact of various configurations of DBudgetKV.
+We select Llama-3-8B-Instruct and perform experiments across five datasets presented in Table 5.
+Appendix B presents additional results and analysis from the ablation study.
+Attention Matrix Reduction
+The reduced attention matrix $A^{\prime}$ in Sec.
+3.2 aggregates attention scores from the last $k$ positions.
+The upper part of Table 5 illustrates the model’s performance and the actual budget when setting $k=1$, $1\%n$, $5\%n$, and $10\%n$ ($n$ being the number of input tokens).
+It is clear that setting $k$ as $1$ achieves a significantly reduced budget, thus a higher compression ratio, with almost no change in performance compared to $1\%n$ and $5\%n$.
+On the other hand, while $10\%n$ can compress more KV cache, it fails to maintain performance (for instance, on NarrativeQA, the former achieves a performance of 9.74 using 32% of the budget, whereas the latter scores 21.44 using 48.7% of the budget).
+What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last token, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length.
+The advantages of both high efficacy and low overhead make $k=1$ a solid design choice for calculating the norm metric.
+Importance Ranking Strategies
+We investigate DBudgetKV potentials through ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O.
+The results are shown in the middle of Table 5: under the same experimental settings, the attention-based ranking struggles to maintain the model’s performance.
+Therefore, the position-based strategy is more suited for our goal.
+Table 6: Performance comparison between DBudgetKV and StreamingLLM with the first two layers frozen (SLMF) under difference KV cache budgets ($0.9$, $0.5$ and $0.2$).
+GSM8K GPQA CoQA NrtvQA QMSum TriviaQA
+Full 75.28 29.02 52.74 24.06 22.27 70.26
+Oursk=1 76.50 30.13 52.86 23.44 21.95 64.24
+Budget 93.2% 86.7% 81.5% 48.7% 15.0% 41.0%
+SLM0.9 72.78 28.79 52.83 23.75 22.48 69.97
+SLMF0.9 76.42 29.24 52.89 22.95 22.47 69.10
+SLM0.5 3.41 6.03 44.68 20.94 21.80 67.79
+SLMF0.5 4.55 3.13 44.59 21.31 21.88 68.04
+SLM0.2 3.21 3.35 40.11 20.12 21.11 66.17
+SLMF0.2 1.21 1.12 38.44 19.52 20.61 65.69
+Attention Norm Threshold
+As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in the bottom part of Table 5.
+Intuitively, a larger threshold allows for a higher compression ratio, while a smaller threshold does the opposite.
+The conclusion is clear that when the threshold is set to a smaller value, such as 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement.
+Conversely, when the threshold is set to 10%, it prunes more KV cache, but the model’s performance significantly deteriorates.
+Thus, we deem 1% as a reasonable threshold for universal uses across models and tasks.
+Comparison with StreamingLLM
+We perform additional experiments on six datasets using StreamingLLM (with Llama3-8B) that has the first two layers frozen, as an ablation study on the effects of the proposed pruning and halting process.
+As shown in Table 6, despite having the first two layers frozen, StreamingLLM still falls short of achieving the objective proposed in this work, not able to remain optimal consistently across datasets like GSM8K and GPQA as the budget varies.
+It is demonstrated that solely leveraging the position bias as in StreamingLLM is not able to maintain the full performance; the proposed pruning and halting mechanism that integrates the attention norm metric is necessary in achieving the goal for lossless pruning.
 5 Conclusion
-In this study, we introduce an innovative KV cache compression objective designed to approximate the full-cache performance, independent of specific inputs, while optimizing resource utilization through targeted KV cache pruning.
+In this study, we introduce a novel KV cache compression objective designed to approximate the full-cache performance, independent of specific inputs, while optimizing resource utilization through targeted KV cache pruning.
 Our approach, termed DBudgetKV, employs a straightforward yet effective two-step process for each layer of LLMs.
@@
 Comprehensive experiments conducted across diverse datasets, encompassing a variety of tasks and context lengths, demonstrate that DBudgetKV achieves nearly lossless compression, while with notable space and time reduction, successfully fulfilling our objective for greater practical values.
-Limitations
-The main limitation of DBudgetKV stems from the gap between its current compression budget and the true optimal budget.
-This can be seen in Table 1 that for certain scenarios, e.g.
-QMSum with Mistral-7B, DBudgetKV reaches 84.3% budget while 50% budget is also viable with no performance degradation.
-We regard this gap as room for improvement in future work, facilitating more aggressive KV cache compression while ensuring the full-cache performance.
-Another limitation of DBudgetKV is that though it demonstrates almost lossless compression, we rely on empirical experiments and there is no hard guarantee on the full-performance constraint.
-In Table 1, while both Llama3-8B and Qwen2.5-7B even surpass the full-cache performance, DBudgetKV with Mistral-7B shows trivial degradation of 1.5%.
-More robust methods could be developed as future work to further strengthen this constraint.
-Appendix A Full Algorithm of DBudgetKV
-Algorithm 1 DBudgetKV
+Appendix A Datasets Used in Experiments
+In this section, we provide a comprehensive overview of all the tasks and datasets utilized in the experiments in this paper.
+Math & Science
+This task evaluates the model’s ability to tackle mathematical and scientific problems.
+By directly inputting questions and comparing the model’s output with the correct answers, we calculate the model’s Accuracy on these datasets: GSM8K is a dataset for evaluating model’s math-solving skills, featuring 8,000 elementary-level math word problems requiring basic arithmetic and reasoning.
+GPQA tests model’s understanding of physics concepts and problem-solving across various topics, assessing scientific reasoning abilities.
+TheoremQA evaluates model’s grasp and application of mathematical theorems, ranging from simple applications to complex proofs, testing advanced math skills.
+Commonsense Reasoning (CR)
+This task evaluates model’s ability to make deductions and understand everyday situations using implicit knowledge and logical inference.
+TruthfulQA (ThQA) evaluates model’s ability to generate accurate and truthful responses, testing models on distinguishing fact from fiction, especially in areas prone to misconceptions.
+We use BLEU as the metric.
+CoQA assesses model’s ability to understand and respond to questions in a conversational context, focusing on maintaining coherence and context throughout a dialogue.
+We use F1 Score as the metric.
+Single Document QA (Single-Doc QA)
+This task assesses the model’s reading comprehension skills when dealing with a single, extended document.
+NarrativeQA [31] is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes.
+Qasper [32] is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information.
+We employ F1 Score as the metric for above two datasets.
+Multi-Document QA (Multi-Doc QA)
+This task evaluates the model’s reading comprehension capabilities across multiple extended documents.
+2WikiMultiHopQA (2WKMQA) [33] is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles.
+MuSiQue [34] evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately.
+We leverage F1 Score as the metric for above two datasets.
+Summarization
+This task examines the model’s ability to comprehend and summarize lengthy documents.
+QMSum [35] is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions.
+Multi-News (M-News) [36] is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic.
+We use Rouge-L as the metric for above two datasets.
+Few-Shot Learning (FSL)
+This task assesses the model’s few-shot learning capabilities.
+TriviaQA [37] is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall.
+We use F1 Score as the metric.
+Code
+This task evaluates the model’s ability to complete and generate code.
+LCC [38] is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures.
+We use Edit Sim as the metric.
+Appendix B Ablation
+Table 7: Performance comparison with different $k$, expanded from Table 5.
+Methods GSM8K CoQA NQA Musique QMSum
+Full 75.28 52.74 24.06 14.77 22.27
+Post=1% Performance Using Different $k$
+$k=1$ 76.50 52.86 23.44 15.96 21.95
+Budget 93.2% 81.5% 48.7% 43.7% 15.0%
+$k=1\%n$ 76.19 52.87 23.17 15.40 21.94
+Budget 95.1% 94.8% 77.3% 77.2% 59.5%
+$k=2\%n$ 76.19 52.82 22.87 14.34 21.74
+Budget 96.0% 96.1% 69.8% 62.0% 47.1%
+$k=3\%n$ 75.82 52.80 23.03 13.61 21.56
+Budget 98.6% 95.7% 60.3% 46.0% 39.5%
+$k=4\%n$ 75.21 52.76 22.96 13.45 21.47
+Budget 98.8% 95.8% 51.9% 35.2% 35.4%
+$k=5\%n$ 75.59 52.75 21.62 13.71 21.43
+Budget 98.6% 96.2% 45.1% 30.0% 33.1%
+$k=6\%n$ 75.66 52.81 21.95 14.33 20.94
+Budget 98.5% 96.6% 40.4% 25.8% 31.6%
+$k=7\%n$ 76.57 52.88 21.18 13.66 20.94
+Budget 98.0% 96.7% 37.3% 22.5% 30.6%
+$k=8\%n$ 76.35 52.86 20.36 13.75 21.11
+Budget 97.4% 96.7% 35.1% 20.9% 30.1%
+$k=9\%n$ 76.65 52.84 19.99 13.66 20.76
+Budget 96.8% 96.8% 33.4% 19.9% 29.5%
+$k=10\%n$ 76.72 52.85 9.74 13.67 21.11
+Budget 96.3% 96.8% 32.0% 19.7% 29.3%
+In this section, we present additional ablation study results for Section 4.5.
+By setting various values for $k$, we expand upon the results shown in Table 5.
+The expanded results are displayed in Table 7.
+These experiments facilitate a deeper understanding of how different parameter settings impact model performance and provide a basis for optimizing parameter selection.
+As shown in Table 7, setting $k=1$ not only conserves pruning time but also achieves better model performance with a reduced budget.
+Appendix C Freeze the first two layers of LLM
+In this section, we demonstrate the necessity of using DBudgetKV to compress the KV cache starting from the 2nd layer of LLMs.
+We apply DBudgetKV on Llama2-7B-Chat and conduct experiments on GSM8K and CoQA.
+We first perform a case study, followed by a comparison of the effects of not freezing layers of the model on its performance.
+As shown in Table 9, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence.
+However, the budget exceeds 90% in both cases.
+By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 0 and 1 are relatively low, while the budgets from layer 2 to layer 30 are the same and very high, with the last layer being low again.
+We analyze that since the model, in the first two layers, is not yet able to identify truly important tokens, the attention distribution is relatively uniform.
+This may lead to the model discarding truly important tokens during eviction, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
+Table 8: Performance on GSM8K and CoQA using DBudgetKV with different frozen layers.
+Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
+GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
+Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
+CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
+Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
+Based on the above case study, we attempt to freeze certain layers of the model and explore the optimal freezing configuration.
+We continue to test the results of freezing different layers.
+The resuls are shown in Table 8.
+From the table, we can observe that freezing the first two layers achieves a balance between model performance and budget.
+Moreover, freezing the 31st layer does not significantly enhance the model’s performance and instead leads to an increase in budget.
+Consequently, DBudgetKV ultimately opts to begin KV cache compression from the 2nd layer of the model.
+Appendix D Full Algorithm of DBudgetKV
   Input: Prompt, Threshold $t$
@@
   return  $K_{c},V_{c}$
-Appendix B Freeze the first two layers of LLM
-In this section, we elaborate the necessity to freeze the first and second layers in DBudgetKV as mentioned in Section 3.
-In our preliminary studies, we apply DBudgetKV to Llama2-7B-Chat and conduct experiments on GSM8K and CoQA.
-We first perform a case study, followed by a comparison with the effects of not freezing the bottom layers.
-As shown in Table 5, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence.
-However, the budget exceeds 90% in both cases.
-By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 0 and 1 are relatively low, while the budgets from layer 2 to layer 30 become very high, with the last layer being low again.
-We hypothesize that the model is still encoding the global semantics in its early layers, thus not yet able to identify truly important tokens in the first two layers, of which the attention distribution is relatively uniform.
-This phenomenon has also been observed by early works on Transformers analysis Ethayarajh (2019); Garí Soler and Apidianaki (2021).
-Not freezing early layers may lead to early eviction of important tokens, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
-Based on the above case study, we attempt to freeze certain layers of the model and explore the optimal freezing configuration.
-We continue to validate the results of freezing different layers, and the resuls are shown in Table 6.
-We can observe that freezing the first two layers achieves a balance between model performance and budget.
-Moreover, freezing the last layer does not significantly enhance the model’s performance and instead leads to an increase in budget.
-DBudgetKV ultimately opts to begin KV cache compression from the 2nd layer of the model.
+Algorithm 1 DBudgetKV
+Table 9: Case Study.
 GSM8K
@@
 Groud-Truth White
-Table 5: Case Study.
-Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
-GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
-Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
-CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
-Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
-Table 6: Performance of Llama2-7B-Chat on GSM8K and CoQA using DBudgetKV with different frozen layers.
-Appendix C Datasets Used in Experiments
-In this section, we provide a comprehensive overview of all the tasks and datasets utilized in the experiments in this paper.
-Math & Science
-This task evaluates the model’s ability to tackle mathematical and scientific problems.
-By directly inputting questions and comparing the model’s output with the correct answers, we calculate the model’s Accuracy on these datasets: GSM8K is a dataset for evaluating model’s math-solving skills, featuring 8,000 elementary-level math word problems requiring basic arithmetic and reasoning.
-GPQA tests model’s understanding of physics concepts and problem-solving across various topics, assessing scientific reasoning abilities.
-TheoremQA evaluates model’s grasp and application of mathematical theorems, ranging from simple applications to complex proofs, testing advanced math skills.
-Commonsense Reasoning (CR)
-This task evaluates model’s ability to make deductions and understand everyday situations using implicit knowledge and logical inference.
-TruthfulQA (ThQA) evaluates model’s ability to generate accurate and truthful responses, testing models on distinguishing fact from fiction, especially in areas prone to misconceptions.
-We use BLEU as the metric.
-CoQA assesses model’s ability to understand and respond to questions in a conversational context, focusing on maintaining coherence and context throughout a dialogue.
-We use F1 Score as the metric.
-Single Document QA (Single-Doc QA)
-This task assesses the model’s reading comprehension skills when dealing with a single, extended document.
-NarrativeQA Kociský et al.
-(2018) is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes.
-Qasper Dasigi et al.
-(2021) is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information.
-We employ F1 Score as the metric for above two datasets.
-Multi-Document QA (Multi-Doc QA)
-This task evaluates the model’s reading comprehension capabilities across multiple extended documents.
-2WikiMultiHopQA (2WKMQA) Ho et al.
-(2020) is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles.
-MuSiQue Trivedi et al.
-(2022) evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately.
-We leverage F1 Score as the metric for above two datasets.
-Methods GSM8K CoQA NQA Musique QMSum
-Full 75.28 52.74 24.06 14.77 22.27
-Post=1% Performance Using Different $k$
-$k=1$ 76.50 52.86 23.44 15.96 21.95
-Budget 93.2% 81.5% 48.7% 43.7% 15.0%
-$k=1\%n$ 76.19 52.87 23.17 15.40 21.94
-Budget 95.1% 94.8% 77.3% 77.2% 59.5%
-$k=2\%n$ 76.19 52.82 22.87 14.34 21.74
-Budget 96.0% 96.1% 69.8% 62.0% 47.1%
-$k=3\%n$ 75.82 52.80 23.03 13.61 21.56
-Budget 98.6% 95.7% 60.3% 46.0% 39.5%
-$k=4\%n$ 75.21 52.76 22.96 13.45 21.47
-Budget 98.8% 95.8% 51.9% 35.2% 35.4%
-$k=5\%n$ 75.59 52.75 21.62 13.71 21.43
-Budget 98.6% 96.2% 45.1% 30.0% 33.1%
-$k=6\%n$ 75.66 52.81 21.95 14.33 20.94
-Budget 98.5% 96.6% 40.4% 25.8% 31.6%
-$k=7\%n$ 76.57 52.88 21.18 13.66 20.94
-Budget 98.0% 96.7% 37.3% 22.5% 30.6%
-$k=8\%n$ 76.35 52.86 20.36 13.75 21.11
-Budget 97.4% 96.7% 35.1% 20.9% 30.1%
-$k=9\%n$ 76.65 52.84 19.99 13.66 20.76
-Budget 96.8% 96.8% 33.4% 19.9% 29.5%
-$k=10\%n$ 76.72 52.85 9.74 13.67 21.11
-Budget 96.3% 96.8% 32.0% 19.7% 29.3%
-Table 7: Performance comparison with different $k$values, ranking methods and universal thresholds.
-Summarization
-This task examines the model’s ability to comprehend and summarize lengthy documents.
-QMSum Zhong et al.
-(2021) is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions.
-Multi-News (M-News) Fabbri et al.
-(2019) is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic.
-We use Rouge-L as the metric for above two datasets.
-Few-Shot Learning (FSL)
-This task assesses the model’s few-shot learning capabilities.
-TriviaQA Joshi et al.
-(2017) is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall.
-We use F1 Score as the metric.
-Code
-This task evaluates the model’s ability to complete and generate code.
-LCC Guo et al.
-(2023) is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures.
-We use Edit Sim as the metric.
-Appendix D Ablation
-In this section, we present additional ablation study results.
-By setting various values for $k$, we expand upon the results shown in Table 3.
-The expanded results are displayed in Table 7.
-These experiments facilitate a deeper understanding of how different parameter settings impact model performance and provide a basis for optimizing parameter selection.
-As shown in Table 7, setting $k=1$ not only conserves pruning time but also achieves better model performance with a reduced budget.
+Appendix E Limitations
+The main limitation of DBudgetKV stems from the gap between its current compression budget and the true optimal budget.
+This can be seen in Table 2 that for certain scenarios, e.g.
+QMSum with Mistral-7B, DBudgetKV reaches 84.3% budget while 50% budget is also viable with no performance degradation.
+We regard this gap as room for improvement in future work, facilitating more aggressive KV cache compression while ensuring the full-cache performance.
+Another limitation of DBudgetKV is that though it demonstrates almost lossless compression, we rely on empirical experiments and there is no hard guarantee on the full-performance constraint.
+In Table 2, while both Llama3-8B and Qwen2.5-7B even surpass the full-cache performance, DBudgetKV with Mistral-7B shows trivial degradation of 1.5%.
+More robust methods could be developed as future work to further strengthen this constraint.
+Appendix F Broader Impacts
+Our work has made contributions on both a new objective in this line of research, as well as a new method that successfully fulfills this objective.
+Since all current KV cache pruning methods require a pre-defined KV budget threshold, they either have inconsistent and unstable performance across different input lengths, domains and tasks, or they need careful tuning of their budgets per dataset to ensure the full-cache performance.
+Our proposed method evaluated on 13 diverse datasets, including short and long-context QA, coding and math, demonstrates its effectiveness on addressing this limitation, which automatically adjusts the budget to reach full-cache performance regardless of the inputs.

### v3

@@
-DBudgetKV: Dynamic Budget in KV Cache Compression for Ensuring Optimal Performance
+Towards Threshold-Free KV Cache Pruning
 Abstract
-To alleviate memory burden during inference of large language models (LLMs), numerous studies have focused on compressing the KV cache by exploring aspects such as attention sparsity.
-These techniques are often designed with a pre-defined KV budget; however, as the optimal budget varies by different input lengths and task types, the existence of a fixed budget could result in inconsistent performance accepting inputs of diverse domains.
-To address this limitation, we propose a new KV cache compression objective: to always ensure the full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible.
-To achieve this goal, we introduce a novel KV cache compression method dubbed DBudgetKV, which features an attention-based metric to signal when the remaining KV cache is unlikely to match the full-cache performance, then halting the pruning process.
-Empirical evaluation spanning diverse context lengths, task types, and model sizes suggests that our method achieves lossless KV pruning effectively and robustly, exceeding 25% compression ratio on average.
-Furthermore, our method is easy to integrate within LLM inference, not only optimizing memory space, but also showing reduced inference time compared to existing methods.
+To reduce memory consumption during LLM inference, prior works have proposed numerous methods that focus on KV cache pruning based on various criteria.
+While these techniques often accomplish lossless memory reduction on many datasets, they often rely on an under-emphasized condition: a dataset/domain-specific budget size threshold needs to be pre-determined to achieve the optimal performance.
+However, such input-specific tuning may be considerably limited in real-world scenarios, as open-domain inputs span diverse domains, lengths and difficulty levels, without clear boundaries for pre-tuning.
+Thus, the dependence of an input-sensitive threshold can be an inherent limitation that may cause large degradation on arbitrary inputs.
+In this work, we propose a new objective that lifts the threshold constraints for robust KV pruning, calling for “threshold-free” methods that automatically adjust budget sizes while ensuring full-cache performance.
+We then propose a novel method ReFreeKV as the first solution fulfilling this objective, validated by intensive experiments on 13 datasets of diverse context lengths, task types, and model sizes.
+Towards Threshold-Free KV Cache Pruning
+Xuanfan Nithanks:  Equal contribution.
+Partial work done during Xuanfan’s internship at WeChat AI.
+Liyan Xu11footnotemark: 1333 Project lead: Liyan Xu <liyanlxu@tencent.com>    Chenyang Lyu   Longyue Wang Mo Yu   Lemao Liu   Fandong Meng   Jie Zhou   Piji Li Nanjing University of Aeronautics and Astronautics WeChat AI, Tencent     Independent Researcher xuanfanni@gmail.com   liyanlxu@tencent.com
 1 Introduction
-Most large language models (LLMs) follows autoregressive generation, where they rely on KV Cache to store intermediate states during inference [1, 2, 3].
-Specifically, LLMs access the cached key and value vectors of past tokens during the attention calculation, thereby speeding up inference by re-using these KV cache [4, 5, 6].
-However, when the model size and the input length increase, the memory required to maintain the KV cache also grows proportionally.
-Since these vectors are usually stored within GPU memory, managing the KV cache efficiently has become crucial to mitigate the overall memory consumption and computational overhead during LLM inference.
-For instance, Llama3 8B model [7] requires 1 GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
-Consequently, recent works have proposed to effectively reduce KV cache, leveraging the sparsity of the attention mechanism: certain positions are more pivotal during the generation process, while those less important ones could be pruned with minimal performance degradation; thus, the full KV cache is not always necessary.
-These observations have led to several optimization methods, such as H2O [8], ScissorHands [9], StreamingLLM [10], and FastGen [11], where they aim to identify and retain only the most salient token positions, discarding less critical ones based on certain pruning criteria.
-Apart from pruning, other paradigms have been proposed as well to reduce KV consumption, such as KVMerger [12] and D2O [13] that merge or compress KV vectors instead of hard-pruning.
-While these studies have successfully demonstrated that a decent portion of KV cache can be discarded in practice, there remains one common issue that has yet to be addressed: these pruning techniques typically require a pre-defined KV cache budget, of which its optimal threshold could vary significantly according to specific tasks or inputs.
-In real-world scenarios, one would need to carefully tune budget thresholds optimized for different domains; otherwise, setting a uniform threshold in diverse domains may potentially bring large degradation on certain tasks.
-We reckon the need for a pre-defined threshold as an overlooked limitation that can lead to unstable and inconsistent performance in practice.
-Table 1: With pre-defined cache budgets, KV pruning methods on Llama3 exhibit inconsistent performance across domains (percentage relative to full-cache scores).
-Methods GSM8K GPQA CoQA
+As current Transformer-based large language models (LLMs) employs autoregressive generation, they produce KV Cache to store intermediate states during inference DBLP:journals/corr/abs-2307-03109; DBLP:journals/corr/abs-2303-08774; DBLP:journals/corr/abs-2402-06196, where they access the cached key and value vectors of past tokens that typically reside within GPU memory for attention calculation DBLP:conf/nips/VaswaniSPUJGKP17; DBLP:journals/corr/abs-1911-02150; DBLP:conf/emnlp/AinslieLJZLS23.
+Consequently, managing the KV cache efficiently has become crucial to mitigate the overall memory consumption and computational overhead during LLM inference, as they can become massive and grow proportionally when the model size and the input length increase.
+For instance, Llama3-8B model DBLP:journals/corr/abs-2407-21783 requires 1GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
+Focusing on this research topic, numerous methods have been recently proposed to effectively reduce the KV footprint after LLM prefilling.
+Leveraging the sparsity of attention, retaining full KV cache is not always necessary.
+Several optimization methods, such as H2O DBLP:conf/nips/Zhang00CZC0TRBW23, ScissorHands DBLP:conf/nips/LiuDLWXXKS23, SnapKV DBLP:conf/nips/LiHYVLYCLC24, FastGen DBLP:conf/iclr/Ge0LZ0024, CAKE DBLP:conf/iclr/QinCLHFCLL25, etc., discard the less critical cache positions based on certain pruning criteria.
+Other paradigms such as KVMerger DBLP:journals/corr/abs-2407-08454 and D2O DBLP:journals/corr/abs-2406-13035 resort to merge or compress KV vectors instead of hard-pruning for achieving the memory reduction effects.
+While these studies have confirmed on a number of datasets that a decent portion of KV cache can be pruned, they hinge on a fundamental yet often under-emphasized condition: a threshold, often being dataset-specific, is typically involved for selective tuning to achieve satisfactory results.
+For instance, D2O takes a pre-defined KV cache budget ratio of 20% to reach full cache performance on LongBench DBLP:conf/acl/BaiLZL0HDLZHDTL24, whereas the required budget jumps to 80% on GSM8K DBLP:journals/corr/abs-2110-14168; likewise, a pre-selected 1024 cache positions is enough for CAKE on LongBench, yet the same budget underperforms on the needle-in-the-haystack test needle.
+The existence of such threshold serves fine in idealized settings given dedicated datasets.
+However, its applicability may be considerably limited in real-world scenarios, where the inputs are intermixed across different domains, lengths and difficulty levels without explicit separation.
+As a result, optimal thresholds cannot be pre-determined for diverse real-world inputs, potentially making the system less robust and prone to significant performance degradation.
+For demonstration, Table 1 presents preliminary experiments using H2O, SnapKV and CAKE where inputs from different datasets are mixed to preclude dataset-specific threshold tuning.
+By varying the KV cache budget ratios, the performance across datasets is shown inconsistent, with a trivial gap under 90% budget but drastically escalating under 20% budget.
+Methods GSM8K GPQA CoQA AVG.
 Budget = 90%
-H2O 99.1% 96.9% 100.2%
-SLM 96.7% 99.2% 100.2%
-SnapKV 93.3% 96.1% 99.9%
+H2O 99.1% 96.9% 100.2% 98.7%
+SnapKV 93.3% 96.1% 99.9% 96.4%
+CAKE 98.6% 93.8% 97.0% 96.5%
 Budget = 20%
-H2O 3.2% 8.5% 96.5%
-SLM 4.3% 11.5% 76.1%
-SnapKV 1.7% 3.9% 98.6%
-To further illustrate, Table 1 shows that when retaining 90% KV cache, Llama3-8B-Instruct could achieve near full-cache performance across three domains.
-However, the performance is shown inconsistent with 20% budget; especially on the math benchmark GSM8K [14], all three KV pruning methods drastically fall to less than 5% relative to the full-cache performance.
-To address the aforementioned limitation, in this work, we propose a new objective for KV cache compression: our method aims to strike towards full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible.
-Towards this objective, we propose an adaptive pruning method, dubbed DBudgetKV, which Dynamically adjusts KV cache retention per input, without the need of pre-defined memory Budget.
-In particular, DBudgetKV automatically identifies input-dependent KV cache pruning that leads to minimal performance degradation.
-A key implication is that the method tends to realize a higher compression ratio for simpler tasks, while allocating more cache resources for complicated tasks.
-Our method features a two-step process: (1) after the prefilling stage, rank all input tokens according to an importance metric; (2) discard the KV cache of each token sequentially, until hitting a stopping criterion.
-In choosing the importance metric and the stopping criterion, DBudgetKV takes into account two important factors from previous works: i) position bias, where tokens at the beginning and the end are generally more significant than those in the middle [10, 15]; ii) attention score, where tokens receiving higher scores usually contribute more information to subsequent generation [1, 2, 8].
-To eliminate the pre-defined KV budget while always targeting full-cache performance, the stopping criterion becomes crucial, such that it should signal whenever the remaining KV cache is unlikely to match the full-cache capacity, thereby fulfilling the objective to approximate full performance with dynamic cache pruning.
-In DBudgetKV, we design a simple attention-based metric as the stopping condition: the metric compares the difference between the Frobenius form of the full attention matrix and the reduced attention matrix of which pruned positions’ scores are set to zero.
-When the difference exceeds a universal threshold, the pruning process halts.
-In practice, we find that this metric empirically correlates well with the full-cache performance regardless of input types, serving as an estimator to bridge the pruning state and the subsequent generation.
-Figure 1: The overall workflow of DBudgetKV in Section 3.2.
-After prefilling, tokens are initially ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix.
+H2O 3.2% 8.5% 96.5% 36.1%
+SnapKV 1.7% 3.9% 98.6% 34.7%
+CAKE 2.2% 15.4% 96.1% 37.9%
+Table 1: KV pruning methods that depend on a preset KV budget threshold can exhibit inconsistent performance across domains (percentage relative to full-cache scores using Llama3), making input-specific threshold selection unavoidable for achieving optimal inference.
+In this work, we propose a new objective that lifts the threshold dependency in KV cache pruning, thereby enabling the system to robustly handle arbitrary inputs without input-specific tuning.
+Specifically, our objective prioritizes two principles: 1) the method operates with a universal threshold insensitive to inputs, effectively rendering it “threshold-free”; and 2) the method should consistently achieve performance comparable to its full-cache counterpart, by automatically adjusting KV cache budgets.
+Only after satisfying these two criteria should the method pursue the best possible compression ratio.
+As shown by our full experimental results in Table 2, while prior methods may achieve strong compression on certain datasets, none satisfies both critical aspects by our new objective.
+Towards our goal, we then propose a novel method, ReFreeKV, featuring a thReshold-Free KV cache pruning.
+Our method adopts a two-step process with a novel metric to dynamically control the KV cache budget.
+Concretely, it first ranks all KV cache positions based on their positional importance, a signal proven useful by StreamingLLM DBLP:conf/iclr/XiaoTCHL24; then, it progressively retains key-value vectors in order, and discards the remaining KV cache once the stopping criterion of our proposed input-insensitive Uni-Metric.
+For minimal overhead, ReFreeKV is implemented with Torch parallel operators, and our analysis in Section 3.3 shows that its latency is on par with prior efficient KV pruning methods varying batch sizes.
+The “threshold-free” aspect stems from the empirical validation of our Uni-Metric described in Section 2.2, whose design is shown to be consistent and insensitive to variations in input domains and sequence lengths.
+Specifically, it adopts a Universal threshold leveraging the Frobenius norm of attention matrix, such that halting by the same threshold empirically incurs minimal degradation relative to the full-cache capacity.
+By design, ReFreeKV naturally yields higher compression ratios for simpler tasks, while allocating more cache resources to more complex ones.
+Figure 1: The overall workflow of ReFreeKV in Section $2.$ After prefilling, tokens are initially ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix.
 The KV cache for the remaining tokens are then preserved to subsequent generation.
-To examine the performance consistency across diverse domains, our experiments adopt 13 datasets varying diverse context lengths and tasks, e.g.
-mathematical and commonsense reasoning, reading comprehension and coding, demonstrating that DBudgetKV achieves our objective effectively and robustly.
-With dynamic KV budgets, the resulting inference is on par or even surpasses the full-cache performance with multiple LLMs of different sizes, including Llama3 [7], Qwen2.5 [16], and Mistral [17].
-The dynamic pruning of DBudgetKV automatically allocates relatively high cache budget on math benchmarks to prevent degradation, while achieving high compression ratio on comprehension tasks by up to 85%.
-With Llama3-8B, DBudgetKV reaches an average budget ratio of 63.7% with minimal degradation across inputs.
-Overall, empirical results suggest that our proposed method is able to accomplish lossless KV cache pruning with strong generalizability.
-Besides, it does not rely on specific model architectures, and is orthogonal to other KV optimization techniques.
-We further conduct experiments for in-depth analysis and ablation studies, where we validate that each step of DBudgetKV is necessary in achieving the desired performance, depicting the rationality of various design choices and hyper-parameters.
-Our contributions can be summarized as follows:
-  • To the best of our knowledge, we are the first to call for the objective that ensures full-cache performance while maximizing the KV cache pruning, addressing the limitation of inconsistent performance with fixed KV budget thresholds.
-  • We introduce DBudgetKV that employs a straightforward two-step process, where the dynamic pruning is halted with minimal degradation to the generation performance.
-  • Extensive evaluation with diverse tasks and models indicates that compared to prior works, our method achieves lossless KV cache compression effectively, efficiently and robustly.
-2 Related Work
-KV Cache Compression
-Recent advancements in KV cache compression have emerged to address memory constraints for model inference.
-Scissorhands [9] observes the repetitive nature of attention patterns and the persistence of token importance.
-It then preserves the most important tokens based on attention scores to achieve KV cache compression.
-H2O [8] introduces dynamic eviction policies that strategically balance retention of recent tokens and historically significant "heavy hitters".
-StreamingLLM [10] enables infinite-length sequence processing without fine-tuning by stabilizing attention through landmark tokens.
-SnapKV [18] enhances compression efficiency through attention score-based selection and clustering of critical KV positions.
-FastGen [11] proposes an adaptive KV cache management system that customizes retention strategies per attention head.
-Other paradigms beyond KV pruning have been proposed along the way, such as KV cache merging [19].
-As prior KV pruning methods typically require setting a pre-defined memory budget, they are by design inherently unable to ensure consistent performance across various domains and datasets, which could prevent their widespread deployment in practice.
-Our method targets to lift such constraint, achieving near full-cache performance and dynamic budget size without the need of additional tuning.
-Attention Pattern
-The study of attention head specialization has paved the way for optimizing memory and computation in Transformer models.
-Early research on encoder architectures shows that some attention heads have persistent positional biases, like consistently focusing on adjacent tokens, indicating redundancy in their learned patterns.
-Rho-1 [20] reveals that the importance of individual heads varies significantly across tasks, suggesting that not all attention computations are equally crucial during inference, and reveal the potential for dynamically pruning or approximating redundant computations.
-Recent works [6, 21] have shifted the focus to decoder-only models, where the quadratic memory overhead of retaining intermediate KV vectors poses a significant challenge.
-Focused Transformer [22] extends the context length of LLMs through a contrastive learning-inspired training process.
-Landmark-Attention [23] uses a “landmark” token to represent each block of the input and trains the attention to use it for selecting relevant blocks, enabling retrieval of blocks directly through the attention mechanism instead of by relying on a separate mechanism.
-Our work leverages attention’s structural regularities, particularly the predictable, position-specific behaviors of certain heads, to dynamically reduce the memory footprint of cached KV states.
-By identifying and exploiting the properties with position bias and attention distribution, we enable efficient inference without compromising model performance, aligning with the broader goal of dynamic KV cache pruning for various inputs.
+To examine its efficacy, our experiments adopt 13 datasets varying diverse context lengths and tasks, e.g.
+mathematical, commonsense reasoning, reading comprehension and coding, demonstrating that ReFreeKV fulfills our objective: without an input-specific threshold, the resulting inference is highly comparable or even surpasses the full-cache performance, evaluated with multiple LLMs of different sizes.
+Using Llama3-8B, it automatically allocates an average KV budget ratio of 63.7%, while retaining 100% full-cache performance across 13 datasets.
+Extensive experiments and ablation studies empirically demonstrate that ReFreeKV effectively accomplishes our proposed objective.
+Moreover, as ReFreeKV does not rely on specific model architectures, it has the potentials to further combine with other orthogonal memory optimization paradigms such as quantization DBLP:conf/nips/HooperKMMSKG24; DBLP:journals/corr/abs-2503-18773; DBLP:journals/corr/abs-2502-04420.
 2 Methodology
-We first elaborate our objective in Sec.
-3.1 as a new direction for KV cache compression.
-Distinct from previous works, we aim to compress the KV cache as much as possible while ensuring LLMs’ performance remains intact.
-We then delineate our proposed approach along with its motivation behind in Sec.
-3.2.
-Important implementation details are next discussed in Sec.
-3.3.
-3.1 Objective of KV Cache Compression
-As mentioned in Sec.
-1, most previous works on KV cache pruning require a pre-defined memory budget beforehand, either by fixing the number of cached positions directly [24, 18], or retaining positions by a fixed ratio of input length [8, 13].
-While these approaches are effective in certain circumstances, they may encounter practical challenges when deployed in real-world scenarios.
-It is forthcoming that the optimal budgets for LLM inference in real-world scenarios are evidently infeasible to enumerate, as they could vary across tasks and domains, especially for open-domain instructions.
-For example, mathematical tasks are typically concise and require inference based on all given conditions, thus many positions may logically contribute to the generation, necessitating a relatively higher memory budget.
-Conversely, for article reading comprehension, LLM may only need a small set of critical positions for the answer generation, featuring a smaller budget.
-As a side effect, inference with fixed KV cache pruning may experience performance instability across different inputs, which can be observed by our analysis in Sec.
-4.2.
-We regard it as a practical limitation that motivates the inception of this work.
-In addressing this limitation, we resort to a dynamic paradigm that eliminates the need for manually setting a memory budget.
-We introduce a new objective for KV cache compression, aiming to automatically reduce the KV cache as much as possible, while fulfilling the condition that the method should always strike towards full-cache performance.
-In this way, one could utilize such pruning method off-the-shelf, without spending time and resources on tuning usage-specific hyper-parameters, which we hope would further advance the research development of KV cache compression techniques.
-3.2 DBudgetKV
-Towards our objective, we design a novel ranking-based pruning method applied after the prefilling stage, comprising two steps.
-First, all tokens are ranked according to an importance metric per Transformers layer.
-Then, the system keeps removing the KV cache sequentially, until hitting a stopping criterion.
-Ultimately, our proposed approach does not need to fix an optimal budget for storing KV cache; rather, it relies on the stopping criterion to determine dynamically, which should work uniformly regardless of the inputs.
-The motivation behind our two-step procedure comes from two phenomena brought by previous works.
-First, there exists position bias such that positions in the beginning and the end are typically important for the subsequent generation, as observed by StreamingLLM [10].
-Second, attention scores are usually an effective indicator of the token importance, as leveraged by many pruning techniques [9, 8, 18].
-For our objective, we propose to combine the two important properties in our approach, described in details as follows.
-Importance Ranking
+In this section, we first elaborate our motivation and the new objective for KV cache compression that differs from prior works.
+We then delineate our proposed approach ReFreeKV, along with its key implementation details.
+2.1 Threshold-Free: A New Objective
+As aforementioned, previous KV cache compression methods require a pre-selected budget threshold.
+While these methods can perform well on numerous datasets, it is forthcoming that the dependence on a threshold can become an inherent limitation.
+As the optimal threshold can vary across different inputs, it may not be feasible for such systems to pick appropriate thresholds or maintain stable performance in practical scenarios involving arbitrary inputs and open-domain instructions.
+Certain inputs may necessitate a relatively higher memory budget, such as tasks with multi-step or mathematical reasoning, whereas certain inputs such as straightforward QA queries need only a small set of KV cache.
+Real-world inputs, however, are mixed and unpredictable, with no clear boundary by difficulty or domain.
+Our propose objective is exactly to remove input-specific threshold constraints, motivating threshold-free methods that ensure consistent performance comparable to full-cache regardless of inputs, obviating the need of tuning for optimal thresholds.
+Pursuit of the best compression ratio is prioritized only after satisfying the objective.
+To the best of our knowledge, we are the first to propose an effective solution that fulfills such objective.
+2.2 ReFreeKV
+ReFreeKV consists of two stages implemented with efficient parallel operators.
+Conceptually, it first ranks all KV cache positions per layer and per attention head.
+It then sequentially retains key–value vectors until a stopping condition, determined by our input-insensitive metric with a universal threshold, is met; the KV cache at the remaining positions is subsequently discarded.
+In line with most prior KV compression works kv_survey, our method is applied once after input prefilling, with the primary goal of reducing memory consumption and an additional benefit of improved throughput (Appendix F).
+Initial Ranking
+The first stage ranks all KV cache, such that the beginning of the sequence may likely contain more critical information than its latter parts, which forms the basis for downstream sequential eviction.
+Building on the positional bias reported in StreamingLLM DBLP:conf/iclr/XiaoTCHL24, we also observe that positions at the beginning and end of the input generally play a more critical role in subsequent generation.
+We exploit this property in the first stage, and rank KV cache by token positions as follows.
 Denote a LLM input sequence as $X=\{x_{1},x_{2},\dots,x_{n}\}$, where each Transformers layer originally consists of $n$ positions of KV vectors per attention head.
-As the pruning process starts from the least important KV vectors, all tokens are firstly ranked according to an importance strategy.
-To this end, we resort to a simple ranking strategy leveraging the position bias, without undergoing any additional computation.
-As identified by StreamingLLM, a significant portion of attentions is allocated to the initial tokens, regardless of their relevance to the language modeling task, known as attention sinks.
-Inspired by this, we employ a purely position-based importance evaluation strategy: we assume that the first $m$ tokens are always crucial; while among the remaining $n-m$ tokens, those towards the end are more significant.
-By this strategy, the importance in $X$ is ranked from the highest to the lowest as $\{x_{1},x_{2},\dots,x_{m},x_{n},x_{n-1},\dots,x_{m+1}\}$, denoted as $\widehat{X}$, where $m$ is a chosen hyperparameter that works regardless of specific input sequences.
-Like previous KV compression works, our choice on importance ranking is empirically supported rather than theoretically based.
-We have also experimented other attention-based importance ranking, which are provided in Sec.
-4.5.
-The choice of position-based ranking is empirically shown both empirically effective and superior on computational overhead.
-Stopping Criterion
-With the initially ranked KV vectors, our method then sequentially decides whether to continue or terminate pruning, which requires an evaluation metric as the stopping criterion.
-Since one evidently cannot know the precise impact of a KV cache position before the generation, we adopt a metric that empirically correlates well with the performance change when discarding a position, serving as a bridge to ensure a minimal degradation upon the full cache performance.
-Inspired by Devoto et al.
-[25], we utilize the Frobenius norm of the attention matrix that fits the requirement well empirically.
-For a matrix $M\in\mathbb{R}^{m\times n}$, the Frobenius norm of $M$ is the matrix-version of the L2 norm, expressed as $||M||_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}|M_{i,j}|^{2}}$.
-Figure 2: Preliminary experiments on the relationship between the Frobenius norm difference (in percentage) obtained from randomly masking tokens, and the corresponding model performance on GSM8K.
-We conduct preliminary experiments to validate its reliability.
-We employ Llama2-7B and randomly mask a certain portion of the KV cache, setting the corresponding values in the attention matrix $A\in\mathbb{R}^{n\times n}$ to zero.
-As illustrated in Figure 2, there is a certain degree of positive correlation between its Frobenius norm and the model’s performance, which are shown generalized across different domains.
-The simple Frobenius norm works empirically well in subsequent large-scale evaluation in Sec.
-4.
-As the input sequence length $n$ increases, the time and space overhead for norm calculation also grows significantly by $O(n^{2})$.
-Hence, we next aim to reduce the computational scale from the naive norm calculation.
-We observe that the Frobenius norm of the last row of $A$ also correlates with the model’s performance; equivalently speaking, the attention distribution of the latest token is good enough to serve as an estimator.
-Formally, we utilize the last $k$ rows (latest $k$ tokens) of $A$ and reduce to a single attention vector $A^{\prime}\in\mathbb{R}^{1\times n}$.
-The score $s_{j}$ for a position $j\in[1,n]$ in $A^{\prime}$ is:
-$s_{j}=\frac{\sum_{i=k}^{n}A_{i,j}}{\sum_{i=k}^{n}\mathbf{1}_{\{A_{i,j}\neq 0\}}}$ (1)
-We then use the Frobenius norm of $A^{\prime}$, equivalently its L2 norm, as the evaluation metric, which reduces the overall complexity to $O(n)$.
-Finally, for each position in the ranking order, DBudgetKV compares the difference between the Frobenius norm of the original attention matrix $A$, and the Frobenius norm of the reduced attention vector $A^{\prime}$ of which the pruned positions are set to 00.
-We set a universal threshold as 1%, such that when the new norm exceeds 1% difference from the original norm, it hits the stopping criterion.
-The KV vectors of the remaining positions are then all kept, while the preceding positions have their KV cache discarded.
-In this way, our method achieves dynamic KV cache pruning while targeting to maintain the full-cache performance, assisted by the norm metric.
-3.3 Implemetation Details
-Though our proposed pruning process is conceptually sequential, it is implemented efficiently by PyTorch’s operators, such that the stopping positions of all layers are identified directly in parallel without any looping operations.
-Empirical analysis in Sec.
-4.4 shows that our proposed method achieves higher efficiency compared to three popular baselines.
-Specifically, we reorder $A^{\prime}$ based on Token Importance Ranking $\widehat{X}$ to obtain $A^{\prime}_{sort}$, where $A^{\prime}_{sort}[0]$ corresponds to the least important token and $A^{\prime}_{sort}[-1]$ corresponds to the most important token.
-Next, we compute the cumulative square-sum of each element in $A^{\prime}_{sort}$:
-$A^{\prime}_{cumsum}[i]=\sqrt{\sum_{k=i}^{n}|A^{\prime}_{sort}[k]|^{2}}$ (2)
-such that $A^{\prime}_{cumsum}[i]$ represents the Frobenius Norm of the attention matrix after removing all tokens to the left of $i^{th}$ token.
-We then divide each element in $A^{\prime}_{cumsum}$ by $||A^{\prime}||_{F}$ and subtract this value from 1 to determine the difference between each new matrix and $A^{\prime}$.
-The torch where operation allows us to directly identify the position of the token that satisfies the condition of not exceeding a 1% difference, ultimately yielding the set of tokens for which the KV cache is called to retain.
-The final kv cache compression algorithm of DBudgetKV is presented in Algorithm 1.
-Since PyTorch allows specifying the dimension for operations, DBudgetKV can run in parallel across multiple attention heads.
+The initial ranking takes the first $m$ positions and reversely takes the remaining nmn-m positions, denoted by $\widehat{X}=\{x_{1},x_{2},\dots,x_{m},x_{n},x_{n-1},\dots,x_{m+1}\}$.
+Note that $m$ is a chosen hyperparameter that works well regardless of specific input sequences.
+Our early experiments also conduct exploration on other ranking strategies, reported in Section 3.4.
+By empirical validation, the position-based ranking is shown not only effective but also particularly advantageous in terms of computational overhead, constituting the first stage for ReFreeKV.
+However, pruning solely by positions does not fulfill our objective, as our experiments in Appendix G show that StreamingLLM can exhibit inconsistent performance varying budget sizes, highlighting the need for a more robust eviction strategy.
+Eviction by Uni-Metric
+With the initial ranking on KV cache, ReFreeKV then sequentially retains KV vectors, and halts upon the stopping condition by an input-insensitive Uni-Metric, after which the remaining cache is effectively evicted.
+The design of Uni-Metric is then at the core of this process, which requires to signal the degradation level after removing KV cache of certain positions.
+we propose a metric that empirically correlates well with the performance change when discarding a position, which could serve as a bridge to ensure a minimal degradation upon the full cache performance.
+Inspired by DBLP:journals/corr/abs-2406-11430, we utilize the Frobenius norm (L2 norm) of the attention matrix $A\in\mathbb{R}^{n\times n}$ as the Uni-Metric, denoted as $A||_{F}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}|A_{i,j}|^{2}}$.
+For each position $i$ in the initially ranked sequence, we compare the Frobenius norm of the original attention matrix, $A||_{F}$, with that of a curated attention matrix, $\widetilde{A_{i}}||_{F}$, in which scores to all positions >i>i in the ranked sequence are masked out, replicating the effect of discarding all KV cache beyond position $i$.
+Once the norm difference reaches a threshold $T$ at position $i_{\text{prune}}$, the entire process terminates, retaining only the KV cache up to $i_{\text{prune}}$ and discarding the remainder, denoted as:
+[…]
+The Universal Threshold
+To fulfill our objective, the threshold $T$ should ensure near lossless pruning invariant to inputs.
+Upon empirical search, we identify $T=1\%$ could serve well for this purpose.
+Figure 2 illustrates how performance varies with changes in the norm across different domains.
+Preliminary studies indicate that when $T<1\%$, performance remains comparable to the full-cache version robustly.
+We select $T=1\%$ to balance minimal degradation with maximal cache eviction.
+The efficacy of Uni-Metric and its universal threshold is validated at full scale in the main experiments presented in Section 3.2 and Figure 3.
+Reducing Overhead
+As the input sequence length $n$ increases, the time and space overhead for norm calculation on the attention matrix grows by $O(n^{2})$.
+To reduce the computational scale, we seek to use an approximate norm calculation by $O(n)$.
+Instead of using the full attention $A$, we reduce $A$ by taking the average of its last $k$ rows to a single attention vector $A^{\prime}\in\mathbb{R}^{1\times n}$.
+The score $s_{i}$ for a position $i\in[1,n]$ in $A^{\prime}$ is denoted as:
+$s_{i}=\frac{\sum_{j=k}^{n}A_{i,j}}{\sum_{j=k}^{n}\mathbf{1}_{\{A_{i,j}\neq 0\}}}$ (2)
+Logistics
+The rationale behind ReFreeKV ’s two-stage process is that finding the optimal pruning positions in a sequence is a combinatorial problem.
+By combining ranking with a linear search, the problem becomes tractable.
+The entire procedure is further optimized through parallel operations, as described as follows.
+Figure 2: Performance trends of Llama3-8B, Mistral-7B, and Qwen2.5-7B across varying Uni-Metric thresholds.
+The x-axis represents the threshold percentage (0.1% to 5%), and the y-axis denotes the performance score normalized by the full-cache performance.
+2.3 Implementation Details
+As ReFreeKV is conceptually a sequential search process, the design of ReFreeKV allows efficient implementation by PyTorch’s operators, such that the stopping positions of all layers are directly identified in parallel without explicit looping operations.
+Latency and throughput analyses in Section 3.3 and Appendix F demonstrate that ReFreeKV matches prior popular KV pruning approaches, with negligible latency overhead and improved throughput compared to the baselines.
+Specifically, the pruning position $i_{\text{prune}}$ can be determined directly by the combination of Torch cumulative-sum and where operators.
+We compute the cumulative square-sum of each element in $A^{\prime}_{\text{rank}}$, obtained after the initial ranking on $A^{\prime}$:
+[…]
+such that $A^{\prime}_{cumsum}[i]$ represents the Frobenius Norm of the attention matrix after removing all cache to the right of position $i$.
+We then divide $\widetilde{A^{\prime}_{i}}||_{F}$ by $A^{\prime}||_{F}$ to determine the norm difference as in Eq (1).
+The torch where operation allows us to directly identify the leftmost position that satisfies the $1\%$ universal threshold, ultimately yielding the set of positions for which the KV cache is called to retain.
+The overall pruning process of ReFreeKV is further presented in Algorithm 1.
 Retaining Bottom Layers
-Additionally, we have discovered that applying DBudgetKV to the model’s 0th and 1st layers results in poor performance when the model operates under a high budget.
-We hypothesize that this may be due to the relatively uniform distribution of attention scores in these initial layers, leading to the premature discarding of tokens that are crucial for subsequent layers, ultimately causing the entire model’s inference to collapse.
-A more detailed analysis of this issue is provided in Appendix C.
-Based on this finding, we initiate the KV cache compression operations starting from the model’s 2nd layer.
-Regarding ablation studies are further conducted in Sec.4.5.
+In our experiments, we discovered that the LLM’s first two Transformers have a relatively uniform attention distribution, usually requiring to retain most of the cache positions.
+For simplicity and robustness, we always retain all KV cache of the first two layers in our implementation.
+More studies on bottom layers are further provided in Appendix B.
+Running at Scale
+Supporting batch sizes >1>1, ReFreeKV is able to perform the entire pruning process for each sample in parallel.
+To achieve this, we pad the shorter cache segments and update the attention masks accordingly, allowing the LLM to ignore the padded KV positions.
+The padding operation has a negligible impact on overall performance.
+Meanwhile, in popular LLM inference engines, e.g.
+vLLM kwon2023efficient, it is possible to allocate separate KV cache size for each sample, which aligns well with ReFreeKV.
+Integration with LLM inference engines such as vLLM is under planning for future development.
 3 Experiments
-4.1 Experimental Settings
-Table 2: Performance of DBudgetKV and its comparison with three KV cache pruning models on 13 datasets.
-Bold numbers indicate the best results aside from full-cache.
-Italics represent the real budget utilized by DBudgetKV.
-The correspondence between abbreviations and their full names of tasks and datasets can be found in Appendix.
-Avg.
-calculates the mean ratio of the model’s performance using different KV cache compression methods to its performance with the full cache.
-All average results (except for budget) are adjusted by subtracting 1 to provide a more intuitive understanding of the effectiveness of different pruning methods.
   Methods Math & Science CR Single-Doc QA Multi-Doc QA Summarization FSL Code   Avg.
@@
 SLM0.9 72.78 28.79 20.75 25.15 52.83 23.75 43.63 32.68 15.86 22.48 27.21 69.97 19.66 -0.59%
-SnapKV0.9 70.20 27.90 20.08 25.38 52.72 24.04 43.93 33.92 15.56 22.49 27.65 70.18 19.05 -1.12%
-H2O0.5 31.08 2.90 16.47 17.99 42.37 23.14 43.05 32.79 15.77 22.79 26.24 69.83 19.18 -17.63%
-SLM0.5 3.41 6.03 8.84 18.75 44.68 20.94 37.73 31.20 15.29 21.80 25.93 67.79 19.12 -24.73%
-SnapKV0.5 12.96 7.81 16.87 19.52 42.84 23.40 43.93 33.98 15.94 22.67 26.07 69.66 19.19 -17.03%
+SnapKV0.9 70.20 27.90 20.08 25.38 52.72 24.04 43.93 33.92 15.56 22.49 27.65 70.58 19.05 -1.12%
+PyramidKV0.9 75.44 28.79 24.46 24.72 52.16 23.64 43.39 32.10 14.11 22.22 23.70 70.48 19.05 -1.59%
+CAKE0.9 74.23 27.23 23.16 22.87 51.18 21.32 43.78 34.65 14.20 21.19 22.33 69.22 19.22 -4.17%
 H2O0.2 2.43 2.46 6.56 20.67 50.91 23.54 42.00 32.50 15.67 22.16 23.68 69.88 19.13 -23.33%
 SLM0.2 3.21 3.35 8.82 18.02 40.11 20.12 36.45 29.91 15.01 21.11 24.68 66.17 18.69 -28.21%
-SnapKV0.2 1.29 1.12 5.89 20.36 52.00 23.10 42.78 32.98 17.01 22.45 24.14 69.63 19.34 -22.89%
+SnapKV0.2 1.29 1.12 5.89 20.36 52.00 23.10 42.78 32.98 14.01 22.45 24.14 69.63 19.34 -22.89%
+PyramidKV0.2 1.36 1.79 4.82 20.24 51.71 24.31 41.90 33.81 12.56 21.11 27.01 68.43 18.19 -25.33%
+CAKE0.2 1.67 4.46 7.93 21.17 50.66 22.01 42.18 33.02 14.93 20.15 24.48 69.92 19.10 -23.47%
 Mistral-7B-Instruct Full 33.36 29.24 6.83 20.82 39.68 28.74 37.80 33.87 22.88 22.19 22.94 86.87 16.08 100%
@@
 SnapKV0.9 27.60 25.67 7.10 20.23 39.76 27.43 38.27 35.66 22.99 22.16 23.31 86.46 16.07 -1.90%
-H2O0.5 2.50 8.26 5.89 20.28 38.53 27.62 36.97 34.31 23.17 22.32 22.60 86.52 16.45 -14.31%
-SLM0.5 2.43 6.47 1.14 18.27 32.76 24.91 33.80 32.73 20.68 21.48 18.54 86.45 13.58 -27.61%
-SnapKV0.5 3.03 8.93 6.83 19.03 39.22 27.15 38.27 34.60 22.94 21.93 22.68 86.31 16.18 -13.41%
+PyramidKV0.9 22.74 26.75 5.35 20.34 38.43 27.60 38.27 37.62 25.28 22.18 24.64 85.53 16.37 -3.15%
+CAKE0.9 25.02 24.10 6.12 20.21 38.82 27.49 39.19 32.96 23.55 22.22 24.51 86.36 15.80 -4.14%
 H2O0.2 1.52 4.46 4.02 18.45 35.55 27.28 33.45 34.36 23.21 21.80 20.93 85.99 14.82 -23.55%
@@
 SnapKV0.2 1.44 0.20 3.35 17.81 37.44 26.76 36.02 33.90 23.05 22.13 20.88 84.27 16.20 -24.28%
+PyramidKV0.2 1.14 0.58 1.40 19.24 31.20 26.19 34.53 37.25 24.55 22.24 22.18 86.58 15.76 -23.75%
+CAKE0.2 4.55 1.03 2.09 18.21 30.07 26.25 36.03 32.39 22.79 21.53 23.96 86.61 15.00 -24.05%
 Qwen2.5-7B-Instruct Full 88.02 31.70 29.85 24.55 61.43 20.81 43.17 47.15 30.70 23.64 24.24 87.64 2.44 100%
@@
 SnapKV0.9 80.14 30.13 28.11 24.30 61.26 20.90 43.31 47.81 30.69 23.90 24.29 87.57 2.55 -1.01%
-H2O0.5 34.12 14.73 20.88 23.10 59.69 21.59 43.37 47.60 30.81 21.00 22.95 85.54 2.24 -13.47%
-SLM0.5 3.26 1.34 10.58 23.75 49.87 19.29 36.45 42.42 25.33 21.24 22.77 87.33 3.40 -23.56%
-SnapKV0.5 20.77 12.28 22.76 22.25 60.21 20.93 43.22 47.45 30.61 23.76 22.70 87.57 2.21 -14.39%
+PyramidKV0.9 84.15 27.68 26.78 23.51 50.34 27.60 42.27 37.62 25.28 22.18 21.65 87.53 3.37 -2.74%
+CAKE0.9 85.67 30.23 29.28 23.52 58.27 29.27 33.73 46.60 30.32 23.91 23.89 90.14 2.41 -0.06%
 H2O0.2 5.91 3.57 17.40 21.70 55.39 21.41 40.30 45.84 29.72 23.04 20.84 87.99 2.15 -13.89%
@@
 SnapKV0.2 2.58 7.37 15.13 21.52 57.06 21.40 40.93 46.96 29.87 23.58 20.69 87.10 2.44 -13.26%
+PyramidKV0.2 9.70 8.57 9.60 26.42 51.19 26.42 39.48 39.54 25.73 21.02 21.65 85.31 1.76 -23.47%
+CAKE0.2 7.58 9.11 5.56 23.66 49.19 30.20 43.04 46.35 29.00 23.33 22.94 85.45 2.89 -16.98%
+Table 2: Performance of ReFreeKV and its comparison with five KV pruning methods on 13 datasets.
+Bold numbers indicate the best results aside from full-cache.
+Italics represent the real budget utilized by ReFreeKV.
+The correspondence between abbreviations and their full names of datasets can be found in Appendix.
+Avg.
+calculates the mean ratio of the model’s performance using different KV cache compression methods to its performance with the full cache.
+All average results (except for budget) are adjusted by subtracting 1 to provide a more intuitive understanding of the effectiveness of different methods.
+3.1 Experimental Settings
 Backbones
-Our experiments are conducted with three LLMs: Llama3-Instruct with size of 8B/70B, Mistral-7B-Instruct-V0.3, and Qwen2.5-Instruct with size of 7B/32B/72B.
-We implement DBudgetKV upon the codebase of SnapKV [^1] https://github.com/FasterDecoding/SnapKV.
-For the reduced attention matrix $A^{\prime}$, we set $k=1$ in practice (ablation provided in Sec.
-4.5), and set $m=4$ for importance ranking.
+Our experiments are conducted with three LLM families of different model sizes: Llama3-Instruct with size of 8B/70B, Mistral-7B-Instruct-V0.3, and Qwen2.5-Instruct with size of 7B/32B/72B.
+We implement our ReFreeKV upon the released codebase of SnapKV***https://github.com/FasterDecoding/SnapKV.
+For the reduced attention matrix $A^{\prime}$, we set $k=1$ in practice (ablation provided in Section 3.4), and set $m=4$ for the initial cache ranking stage.
 Datasets
-For comprehensive evaluation, we evaluate DBudgetKV using datasets of both short and long context length.
-For tasks of relatively short inputs, the models are assessed across mathematics, science, and commonsense reasoning on multiple datasets including GSM8K [14], GPQA [26], TheoremQA [27], TruthfulQA [28], and CoQA [29].
-For long-context tasks, we adopt tasks and datasets from Longbench [30].
-Appendix A provides a detailed description and statistics of these datasets along with how they are utilized in experiments.
+For comprehensive evaluation, we evaluate ReFreeKV on datasets of both short and long context length of different domains, including mathematics, science, and commonsense reasoning on GSM8K DBLP:journals/corr/abs-2110-14168, GPQA DBLP:journals/corr/abs-2311-12022, TheoremQA DBLP:conf/emnlp/ChenYKLWMXWX23, TruthfulQA DBLP:conf/acl/LinHE22, and CoQA DBLP:journals/tacl/ReddyCM19.
+We also include tasks from Longbench DBLP:conf/acl/BaiLZL0HDLZHDTL24 with 8 datasets spanning document comprehension, summarization and coding.
+Appendix D provides a detailed description and statistics for all 13 datasets, along with how they are utilized in our experiments.
 Evaluation Protocol
-Our primary objective is to evaluate the capability of DBudgetKV to compress KV cache while striking for a full-cache performance.
-To this end, we compare DBudgetKV with the full KV cache performance and calculate the average compression ratio per dataset.
-Additionally, we select three previous KV cache compression methods with fixed budget size, including: Heavy Hitter Oracle (H2O) [8], StreamingLLM (SLM) [10], and SnapKV [18].
-We compare with three budgets for these three methods at 90%, 50% and 20% compression ratio respectively.
-Apart from the main experiments, we further evaluate various design factors of DBudgetKV via ablation studies, such as different strategies of token ranking importance, the selection of $k$ for the attention matrix, and the necessity to retain bottom layers.
-Finally, we also provide efficiency analysis on the time and space overhead comparing against other KV cache pruning methods in Sec.
-4.4.
+Our evaluation primarily assesses ReFreeKV ’s ability to preserve full-cache performance with its automatic pruning budgets.
+Accordingly, we compare ReFreeKV against its full-cache scores and also report the average compression ratio for each dataset.
+Additionally, we compare with five prior KV cache pruning methods with varied budget sizes, including: Heavy Hitter Oracle (H2O) DBLP:conf/nips/Zhang00CZC0TRBW23, StreamingLLM (SLM) DBLP:conf/iclr/XiaoTCHL24, SnapKV DBLP:conf/nips/LiHYVLYCLC24, PyramidKV DBLP:journals/corr/abs-2406-02069 and CAKE DBLP:conf/iclr/QinCLHFCLL25.
+By evaluating performance consistency under fixed KV cache budgets of 90%, 50%, and 20%, we further illustrate the limitations arising from the input-specific threshold dependence.
+Lastly, we also include a concurrent work Twilight DBLP:journals/corr/abs-2502-02770, which does not rely on a budget threshold but employs a top-$p$-inspired metric for adaptive token selection.
+It is worth noting that the $p$ value remains a hyperparameter to be tuned across models and inputs.
 3.2 Main Results
-The main results are shown in Table 2 according to our evaluation protocol, where DBudgetKV demonstrates strong advantages with robust performance across diverse tasks and models:
-Table 3: Performance of LLMs with scales of 70B, 32B, and 72B across five datasets using DBudgetKV.
-GSM8K CoQA NQA Musique QMSum
-Llama3-70B-Instruct
-Full 89.69 60.36 27.15 29.31 22.52
-Oursk=1 89.76 60.38 26.94 28.88 22.30
-Budget 93.3% 77.2% 58.7% 60.8% 52.9%
-Qwen2.5-32B-Instruct
-Full 91.36 58.03 24.78 40.04 22.84
-Ours 91.81 57.29 22.65 40.54 22.52
-Budget 91.6% 85.0% 68.2% 74.4% 76.8%
-Qwen2.5-72B-Instruct
-Full 90.22 54.14 24.36 42.13 23.93
-Ours 90.30 54.19 24.10 41.70 23.31
-Budget 90.9% 87.6% 63.4% 65.0% 68.6%
-• DBudgetKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types.
-Interestingly, with Llama3-8B and Qwen2.5-7B, DBudgetKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache.
-With Mistral, DBudgetKV also manages to achieve nearly 20% KV cache compression with only a 1.5% performance reduction.
-These results indicate that our proposed method can be suited for real-world scenarios with no bother by domain-specific budget thresholds.
+The main experimental results are shown in Table 2.
+Due to space limitations, the results for the 50% budget size are placed in Appendix C.
+For comparison with Twilight, we separately report the results in Table 4 due to its different hyperparameter type.
+Based on Table 2, we can draw the following observations.
+• ReFreeKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types.
+Interestingly, with Llama3-8B and Qwen2.5-7B, ReFreeKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache.
+With Mistral, ReFreeKV also manages to achieve near 15% compression with a relatively small 1.5% performance reduction.
+These results indicate that our proposed method can be suited for real-world scenarios with no bother by input-specific budget thresholds.
 • In stark contrast, previous methods achieve a consistent full-cache performance only when manually determined a high budget ratio, e.g.
@@
 However, when the budget is reduced, e.g.
-50% and 20%, the degradation can become severe on certain datasets, distinct from DBudgetKV that automatically adjusts the pruning to always prioritize full performance.
-• DBudgetKV naturally reflects the difficulty of the generation task.
+50% and 20%, the degradation can become severe on certain datasets, distinct from ReFreeKV that automatically adjusts the pruning to always prioritize full-cache performance.
+Theoretically, one could tune the budget for each dataset that achieves minimal degradation, but this is generally infeasible for real-world open-domain instructions.
+• ReFreeKV naturally reflects the difficulty of the generation task.
 As in Table 2, the dynamic budget ratio is high on Math&Science datasets (over 90%), while much lower on QA or Summarization datasets (as low as 15%).
 This observation is in line with our intuition, where inference on concise but hard tasks, such as math problems, requires more context and more precise calculation, resulting in higher budget allocation.
-Thus, our method design serves beyond for inference efficiency, but could also be leveraged for input analysis in a broader scope.
-• Besides the dynamic compression, DBudgetKV also outperforms the three 90%-budget baselines, while itself uses less than 90% budget, and achieves the highest score in 20 out of 39 comparisons.
-This suggests that by taking into account both the position bias and attention scores, our method accompolishes more effective pruning than the baselines that only consider one aspect, highlighting the importance of integrating both dimensions for effective KV cache compression.
-It is worth noting that the scope of this work is not to propose yet another KV cache pruning method.
-Instead, we identify the common limitation of previous KV pruning methods, and call for optimal KV pruning without a pre-defined budget.
-Consequently, the consistent performance across domains after KV pruning is the ultimate goal of our method, instead of achieving the highest compression ratio on certain datasets.
-Under our objective, we leave the further improvement on compression ratio as a promising future research direction.
-4.3 DBudgetKV’s Generalization Ability
-We next conduct experiments to examine whether our design and hyper-parameters can be generalized to LLMs of larger sizes.
-On five datasets in Table 3, DBudgetKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same DBudgetKV setting.
-Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% compression.
-But most importantly, they are able to maintain lossless performance across inputs, validating the robustness and versatility for its general applicability.
-4.4 Efficiency Analysis
-Table 4: The average inference time and pruning time of Llama3 with size of 8B/70B across six datasets, measured in seconds, with lower values indicating better performance.
+From this aspect, our method design serves beyond for memory efficiency, but could be potentially leveraged for input analysis in a broader scope.
+• Besides the dynamic compression, ReFreeKV also outperforms the three 90%-budget baselines, while itself uses less than 90% budget.
+Though, It is worth noting that the goal of this work is not to propose yet another KV cache pruning method that targets the best possible compression ratio under specific conditions.
+Instead, we seek to lift the threshold constraints and advocate for robust KV pruning that generalizes to arbitrary inputs.
+Through our proposed objective, we hope to further advance research in this direction.
 GSM8K CoQA NarrativeQA Musique QMSum TriviaQA Avg.
@@
 SnapKV0.5 16.517 0.117 1.116 0.241 6.795 2.474 7.085 2.468 13.846 2.793 5.562 1.083 8.487 1.529
-Table 5: Performance comparison for ablation studies in Section 4.5, with different $k$ values for attention matrix reduction, ranking methods and universal thresholds.
+Table 3: The average inference time and pruning time of Llama3 with size of 8B/70B across six datasets, measured in seconds, with lower values indicating better performance.
+Methods GSM8K NQA Qasper 2WQA Musique
+Llama3-8B 75.28 24.06 43.91 35.33 14.77
++ Oursk=1 76.50 23.44 37.38 36.21 15.96
++ Twip=0.95 76.04 23.37 43.08 36.18 14.92
+Mistral-7B 33.36 28.74 37.80 33.87 22.88
++ Oursk=1 31.54 27.20 38.93 33.46 22.15
++ Twip=0.85 30.33 27.17 38.90 33.23 22.92
+Table 4: Performance comparison between ReFreeKV and Twilight across five datasets using Llama3-8B-Instruct and Mistral-7B-V0.3. Bold numbers indicate the best results aside from full-cache.
+As separately reported in Table 4, we compare ReFreeKV with Twilight’s reported performance across the GSM8K, NarrativeQA, 2WikiMQA, and Musique datasets.
+Regarding the hyperparameter for Twilight, we adopt their reported optimal $p$ value: $p=0.95$ for Llama3-8B and $p=0.85$ for Mistral.
+As shown in the results, ReFreeKV achieves performance comparable to Twilight on both models.
+Notably, both methods frequently surpass the full-cache baseline across these datasets.
+Though, as Twilight requires different optimal $p$ values for each model, tuning may still be necessary.
+Nonetheless, both ReFreeKV and Twilight share the same principle of adaptive pruning, which we believe can further advance research in this area.
+3.3 Efficiency Analysis
+In this section, we conduct a quantitative study on the latency of ReFreeKV and the overall impact on the inference time, apart from the space reduction from KV cache pruning.
+We compare the time usage on six datasets with Llama3-8B/70B using ReFreeKV, along with three baselines with the 50% budget setting.
+Table 3 reports the time latency needed to apply pruning (Prune), as well as the average generation time after the prefilling stage of each sample (Overall).
+From the results, the latency by ReFreeKV is on par with prior pruning methods.
+Notably, ReFreeKV achieves the best performance in 8 out of 12 comparisons of overall generation time, suggesting that the total generation speed of ReFreeKV is advantageous.
+The trend also remains consistent across different model sizes, underscoring the efficiency aspect of ReFreeKV.
+For batch process, we further conduct analysis on throughput with different batch sizes.
+The full details of which are provided in Appendix F.
+The results confirm that ReFreeKV improves throughput upon the naive generation by 10-20%, especially it retains performance edges and robustness with increasing batch sizes.
+3.4 Ablation Studies
+We conduct ablation studies to investigate the impact of various configurations of ReFreeKV.
+We use Llama-3-8B-Instruct and perform experiments across five datasets, with results presented in Table 5.
+Appendix E provides additional results and analysis from our ablation studies.
+Figure 3: Performance vs.
+Efficiency Trade-off.
+(a) Performance retention across five datasets (solid lines).
+(b) Computational budget consumption (dashed lines) relative to the dense baseline.
+The shared legend indicates the datasets.
+Results show that setting the universal threshold to $1\%$ could well balance between performance and memory budget, as it maintains robust full-cache performance while substantially reducing KV cache.
 Methods GSM8K CoQA NQA Musique QMSum
@@
 Budget 61.7% 28.6% 6.7% 7.9% 7.1%
-Posk=1 Performance with Different Threshold $t$
-${t=0.1\%}$ 76.35 52.84 22.73 17.04 21.90
-Budget 99.2% 97.0% 77.2% 85.0% 73.9%
-${t=10\%}$ 26.23 49.69 18.35 15.16 18.81
-Budget 55.8% 37.2% 15.27% 18.5% 6.5%
-In this section, a quantitative study is conducted on whether DBudgetKV improves inference time, apart from the space reduction from KV cache pruning.
-We compare the time usage on six datasets with Llama3-8B/70B using DBudgetKV, along with three baselines with the 50% budget setting.
-Table 4 reports the average generation time after the prefilling stage of each sample (Overall), as well as the time needed to complete pruning (Prune) before the generation.
-From the results, DBudgetKV achieves the best performance in 8 out of 12 comparisons of overall generation time, and achieves the best average time for both 8B and 70B LLMs, suggesting that although DBudgetKV requires a slightly longer time for pruning itself compared to three other approaches, the total generation speed of DBudgetKV is evidently advantageous.
-It also performs consistently across models of different scales, underscoring its time efficiency.
-4.5 Ablation Study
-The ablation study is conducted to explore the impact of various configurations of DBudgetKV.
-We select Llama-3-8B-Instruct and perform experiments across five datasets presented in Table 5.
-Appendix B presents additional results and analysis from the ablation study.
+Table 5: Performance comparison for ablation studies in Section 3.4, with different $k$ values for attention matrix reduction and initial ranking alternatives.
 Attention Matrix Reduction
-The reduced attention matrix $A^{\prime}$ in Sec.
-3.2 aggregates attention scores from the last $k$ positions.
+The reduced attention matrix $A^{\prime}$ in Section 2.2 aggregates attention scores from the last $k$ rows.
 The upper part of Table 5 illustrates the model’s performance and the actual budget when setting $k=1$, $1\%n$, $5\%n$, and $10\%n$ ($n$ being the number of input tokens).
@@
 On the other hand, while $10\%n$ can compress more KV cache, it fails to maintain performance (for instance, on NarrativeQA, the former achieves a performance of 9.74 using 32% of the budget, whereas the latter scores 21.44 using 48.7% of the budget).
-What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last token, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length.
+What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last row, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length.
 The advantages of both high efficacy and low overhead make $k=1$ a solid design choice for calculating the norm metric.
-Importance Ranking Strategies
-We investigate DBudgetKV potentials through ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O.
-The results are shown in the middle of Table 5: under the same experimental settings, the attention-based ranking struggles to maintain the model’s performance.
-Therefore, the position-based strategy is more suited for our goal.
-Table 6: Performance comparison between DBudgetKV and StreamingLLM with the first two layers frozen (SLMF) under difference KV cache budgets ($0.9$, $0.5$ and $0.2$).
-GSM8K GPQA CoQA NrtvQA QMSum TriviaQA
-Full 75.28 29.02 52.74 24.06 22.27 70.26
-Oursk=1 76.50 30.13 52.86 23.44 21.95 64.24
-Budget 93.2% 86.7% 81.5% 48.7% 15.0% 41.0%
-SLM0.9 72.78 28.79 52.83 23.75 22.48 69.97
-SLMF0.9 76.42 29.24 52.89 22.95 22.47 69.10
-SLM0.5 3.41 6.03 44.68 20.94 21.80 67.79
-SLMF0.5 4.55 3.13 44.59 21.31 21.88 68.04
-SLM0.2 3.21 3.35 40.11 20.12 21.11 66.17
-SLMF0.2 1.21 1.12 38.44 19.52 20.61 65.69
-Attention Norm Threshold
-As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in the bottom part of Table 5.
-Intuitively, a larger threshold allows for a higher compression ratio, while a smaller threshold does the opposite.
-The conclusion is clear that when the threshold is set to a smaller value, such as 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement.
+Initial Ranking Strategies
+Apart from the position-based ranking described in Section 2.2, we also investigate other alternatives, such as ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O.
+The results are shown in the bottom of Table 5: under the same experimental settings, the attention-based ranking struggles to maintain a robust final performance.
+Empirical validation supports that position ranking is an appealing choice, offering both superior efficacy and efficiency.
+The Universal Threshold
+As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in Figure 3.
+Intuitively, a larger threshold allows for a higher compression ratio but potentially more degradation, while a smaller threshold does the opposite.
+It is clear that when the threshold is set to a smaller value, such as 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement.
 Conversely, when the threshold is set to 10%, it prunes more KV cache, but the model’s performance significantly deteriorates.
-Thus, we deem 1% as a reasonable threshold for universal uses across models and tasks.
-Comparison with StreamingLLM
-We perform additional experiments on six datasets using StreamingLLM (with Llama3-8B) that has the first two layers frozen, as an ablation study on the effects of the proposed pruning and halting process.
-As shown in Table 6, despite having the first two layers frozen, StreamingLLM still falls short of achieving the objective proposed in this work, not able to remain optimal consistently across datasets like GSM8K and GPQA as the budget varies.
-It is demonstrated that solely leveraging the position bias as in StreamingLLM is not able to maintain the full performance; the proposed pruning and halting mechanism that integrates the attention norm metric is necessary in achieving the goal for lossless pruning.
+Thus, we deem 1% as a reasonable and robust threshold for universal uses across models and tasks.
+Generalization
+We further conduct experiments to examine whether our design and hyperparameters can be generalized to LLMs of more sizes.
+On five datasets in Appendix Table 12, ReFreeKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same ReFreeKV setting.
+Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% compression.
 4 Conclusion
-In this study, we introduce a novel KV cache compression objective designed to approximate the full-cache performance, independent of specific inputs, while optimizing resource utilization through targeted KV cache pruning.
-Our approach, termed DBudgetKV, employs a straightforward yet effective two-step process for each layer of LLMs.
-Tokens are initially ranked solely based on their positions, followed by the eviction of the least significant tokens as determined by the norm value of the reduced attention matrix.
-Comprehensive experiments conducted across diverse datasets, encompassing a variety of tasks and context lengths, demonstrate that DBudgetKV achieves nearly lossless compression, while with notable space and time reduction, successfully fulfilling our objective for greater practical values.
-Appendix A Datasets Used in Experiments
-In this section, we provide a comprehensive overview of all the tasks and datasets utilized in the experiments in this paper.
-Math & Science
-This task evaluates the model’s ability to tackle mathematical and scientific problems.
-By directly inputting questions and comparing the model’s output with the correct answers, we calculate the model’s Accuracy on these datasets: GSM8K is a dataset for evaluating model’s math-solving skills, featuring 8,000 elementary-level math word problems requiring basic arithmetic and reasoning.
-GPQA tests model’s understanding of physics concepts and problem-solving across various topics, assessing scientific reasoning abilities.
-TheoremQA evaluates model’s grasp and application of mathematical theorems, ranging from simple applications to complex proofs, testing advanced math skills.
-Commonsense Reasoning (CR)
-This task evaluates model’s ability to make deductions and understand everyday situations using implicit knowledge and logical inference.
-TruthfulQA (ThQA) evaluates model’s ability to generate accurate and truthful responses, testing models on distinguishing fact from fiction, especially in areas prone to misconceptions.
-We use BLEU as the metric.
-CoQA assesses model’s ability to understand and respond to questions in a conversational context, focusing on maintaining coherence and context throughout a dialogue.
-We use F1 Score as the metric.
-Single Document QA (Single-Doc QA)
-This task assesses the model’s reading comprehension skills when dealing with a single, extended document.
-NarrativeQA [31] is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes.
-Qasper [32] is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information.
-We employ F1 Score as the metric for above two datasets.
-Multi-Document QA (Multi-Doc QA)
-This task evaluates the model’s reading comprehension capabilities across multiple extended documents.
-2WikiMultiHopQA (2WKMQA) [33] is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles.
-MuSiQue [34] evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately.
-We leverage F1 Score as the metric for above two datasets.
-Summarization
-This task examines the model’s ability to comprehend and summarize lengthy documents.
-QMSum [35] is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions.
-Multi-News (M-News) [36] is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic.
-We use Rouge-L as the metric for above two datasets.
-Few-Shot Learning (FSL)
-This task assesses the model’s few-shot learning capabilities.
-TriviaQA [37] is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall.
-We use F1 Score as the metric.
-Code
-This task evaluates the model’s ability to complete and generate code.
-LCC [38] is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures.
-We use Edit Sim as the metric.
-Appendix B Ablation
-Table 7: Performance comparison with different $k$, expanded from Table 5.
+In this study, we introduce a new KV cache compression objective that lifts the threshold dependency, so to achieve input-insensitive pruning for robust inference performance.
+Towards this objective, we propose a novel method, termed ReFreeKV, which employs a straightforward yet effective two-stage KV cache pruning process.
+Comprehensive experiments conducted across diverse datasets, encompassing a variety of tasks and context lengths, demonstrate that ReFreeKV achieves nearly lossless compression without input-specific tuning or thresholds, while with notable memory reduction.
+Limitations
+The main limitation of ReFreeKV stems from the gap between its current compression budget and the true optimal budget.
+This can be seen in Table 2 that for certain scenarios, e.g.
+QMSum with Mistral-7B, ReFreeKV reaches 84.3% budget while 50% budget is also viable with no performance degradation.
+We regard this gap as room for improvement in future work, facilitating more aggressive KV cache compression while ensuring the full-cache objective.
+Another limitation of ReFreeKV is that though it demonstrates almost lossless compression, we rely on empirical experiments, and there is no hard guarantee on the degree of degradation.
+In Table 2, while both Llama3-8B and Qwen2.5-7B even surpass the full-cache performance, ReFreeKV with Mistral-7B shows trivial degradation of 1.5%.
+More robust methods could be developed as future work to further strengthen the robustness.
+Appendix A Full Algorithm of ReFreeKV
+Algorithm 1 ReFreeKV
+Input: Prompt, Threshold $t$
+Output: Compressed KV Cache
+ Create Empty List $K_{c},V_{c}$
+for Transformer Layer $L_{i}$ in LLM do
+  $Q^{i},K^{i},V^{i}\leftarrow L_{i}(\text{Prompt})$
+  $R^{i}\leftarrow$ Postion-Based Importance Rank
+  $A_{last}^{i}\leftarrow\operatorname{Attention}(Q^{i}[\dots,-1,:],K^{iT})$
+  $F_{b}^{i}\leftarrow\operatorname{Frobenius}(A_{last}^{i})$
+  $A_{last}^{i}\leftarrow\operatorname{Square}(A_{last}^{i})$
+  Reorder $A_{last}^{i}$ by Rank $R$
+  $A_{cumsum}^{i}\leftarrow\operatorname{Cumsum}(A_{last}^{i})$
+  $A_{cumsum}^{i}\leftarrow\operatorname{Sqrt}(A_{cumsum}^{i})$
+  $A_{ratio}^{i}\leftarrow(F_{b}^{i}-A_{cumsum}^{i})/F_{b}^{i}$
+  $\text{Index}^{i}\leftarrow\operatorname{Max}(\operatorname{Where}(A_{ratio}^{i}<=t))$
+  $K_{c}^{i}\leftarrow\operatorname{Compress}K^{i}\text{by}R^{i}[I:]$
+  $V_{c}^{i}\leftarrow\operatorname{Compress}V^{i}\text{by}R^{i}[I:]$
+  Append $K_{c}^{i},V_{c}^{i}$ to $K_{c},V_{c}$
+end for
+return $K_{c},V_{c}$
+Appendix B Retain the first two layers of LLM
+In this section, we demonstrate the design choice of ReFreeKV to compress the KV cache starting from the third layer of LLMs.
+We apply ReFreeKV with Llama2-7B-Chat and conduct experiments on GSM8K and CoQA.
+We first perform a case study, followed by a comparison of the effects of not retaining full KV cache of the first two layers.
+As shown in Table 6, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence.
+However, the budget exceeds 90% in both cases.
+By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 1 and 2 are relatively low, while the budgets from layer 3 to layer 31 are the same and very high, with the last layer being low again.
+We analyze that since the model, in the first two layers, is not yet able to identify truly important tokens, the attention distribution is relatively uniform.
+This may lead to the model discarding truly important tokens during eviction, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
+Based on the above case study, we attempt to retain certain layers of the model and explore the optimal configuration.
+We continue to test the results of retain different layers.
+The results are shown in Table 7.
+We can observe that freezing the first two layers achieves a balance between model performance and budget.
+Moreover, retaining the 31st layer does not significantly enhance the model’s performance and instead leads to an increase in budget.
+Consequently, ReFreeKV ultimately opts to begin KV cache compression from the 3rd layer of the model.
+GSM8K
+Input
+A robe takes 2 bolts of blue fiber a-
+nd half that much white fiber. How
+many bolts in total does it take?
+Budget
+Layer 1: 67.71
+Layer 2: 82.29
+Layer $\sim3$1: 95.83
+Layer 32: 37.50
+Avg.: 93.90
+Output
+I have determined by answering the
+answer to the format of bolts bolts b-
+olts…(repeat)
+Ground-Truth 3
+CoQA
+Input
+You are given a story and a question.
+Answer the question as concisely as
+you can…Question: What color was
+Cotton?
+Budget
+Layer 1: 42.14
+Layer 2: 42.54
+Layer $\sim3$1: 99.19
+Layer 32: 79.64
+Avg.: 95.03
+Output
+Question: Question: What is the que-
+stion: What is the question:
+Groud-Truth White
+Table 6: Case Study.
+Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
+GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
+Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
+CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
+Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
+Table 7: Performance of Llama2-7B-Chat on GSM8K and CoQA using ReFreeKV with different frozen layers.
+Appendix C Main Results with 50% Budget
+  Methods Math & Science CR Single-Doc QA Multi-Doc QA Summarization FSL Code   Avg.
+GSM8K
+GPQA
+TheoQA
+ThQA
+CoQA
+NrtvQA
+Qasper
+2WkMQA
+Musique
+QMSum
+M-News
+TriviaQA
+Lcc
+Llama3 Full 75.28 29.02 21.29 25.59 52.74 24.06 43.91 35.33 14.77 22.27 27.37 70.26 19.16 100%
+Oursk=1 76.50 30.13 23.03 26.09 52.86 23.44 37.38 36.21 15.96 21.95 27.88 64.24 19.31 +0.12%
+Budget 93.2% 86.7% 92.8% 78.7% 81.5% 48.7% 46.4% 45.8% 43.7% 15.0% 76.4% 41.0% 78.0% 63.68%
+H2O0.5 31.08 2.90 16.47 17.99 42.37 23.14 43.05 32.79 15.77 22.79 26.24 69.83 19.18 -17.63%
+SLM0.5 3.41 6.03 8.84 18.75 44.68 20.94 37.73 31.20 15.29 21.80 25.93 67.79 19.12 -24.73%
+SnapKV0.5 12.96 7.81 16.87 19.52 42.84 23.40 43.93 33.98 15.94 22.67 26.07 69.66 19.19 -17.03%
+Mistral Full 33.36 29.24 6.83 20.82 39.68 28.74 37.80 33.87 22.88 22.19 22.94 86.87 16.08 100%
+Oursk=1 31.54 29.02 6.71 20.81 39.80 27.20 38.93 33.46 22.15 22.39 22.67 86.81 15.33 -1.50%
+Budget 89.6% 90.5% 84.2% 92.3% 84.1% 78.0% 97.9% 86.7% 74.4% 84.3% 89.1% 87.2% 89.4% 86.75%
+H2O0.5 2.50 8.26 5.89 20.28 38.53 27.62 36.97 34.31 23.17 22.32 22.60 86.52 16.45 -14.31%
+SLM0.5 2.43 6.47 1.14 18.27 32.76 24.91 33.80 32.73 20.68 21.48 18.54 86.45 13.58 -27.61%
+SnapKV0.5 3.03 8.93 6.83 19.03 39.22 27.15 38.27 34.60 22.94 21.93 22.68 86.31 16.18 -13.41%
+Qwen2.5 Full 88.02 31.70 29.85 24.55 61.43 20.81 43.17 47.15 30.70 23.64 24.24 87.64 2.44 100%
+Oursk=1 88.02 31.25 30.39 24.48 60.01 20.66 42.74 47.20 29.56 22.64 24.04 87.65 3.58 +2.63%
+Budget 90.7% 88.8% 86.5% 69.2% 84.1% 65.1% 69.2% 73.3% 64.4% 70.6% 85.4% 56.6% 84.4% 76.02%
+H2O0.5 34.12 14.73 20.88 23.10 59.69 21.59 43.37 47.60 30.81 21.00 22.95 85.54 2.24 -13.47%
+SLM0.5 3.26 1.34 10.58 23.75 49.87 19.29 36.45 42.42 25.33 21.24 22.77 87.33 3.40 -23.56%
+SnapKV0.5 20.77 12.28 22.76 22.25 60.21 20.93 43.22 47.45 30.61 23.76 22.70 87.57 2.21 -14.39%
+Table 8: Performance of ReFreeKV and its comparison with five KV cache pruning models on 13 datasets. The cache size is set to 50%.
+We show peformance comparison between ReFreeKV and different eviction-based methods with 50% cache size.
+The results are show in Table 8.
+From the table, we can observe the same conclusions with Section 3.
 Methods GSM8K CoQA NQA Musique QMSum
@@
 Budget 96.3% 96.8% 32.0% 19.7% 29.3%
+Table 9: Performance comparison with different $k$, expanded from Table 5.
+BS=1 BS=2 BS=8 BS=16
+[Input, Output] [4K, 8K] [8K, 16K] [512, 4K] [4K, 8K] [512,512] [4K, 4K] [512, 512] [2K, 2K]
+Latency (s/100tokens)
+HF Accelerate 3.61 5.29 3.50 5.26 3.48 12.23 4.71 13.01
+Ours 3.19 4.08 2.84 3.60 3.06 10.42 4.25 10.64
+Throughput (token/s)
+HF Accelerate 27.71 18.89 57.08 38.01 230.00 65.41 339.43 122.98
+ReFreeKV 31.35 24.51 70.30 55.56 261.83 76.81 376.89 150.40
+Budget 73.3% 55.9% 62.0% 34.0% 76.7% 65.3% 82.8% 78.2%
+Table 10: Performance comparison across various batch sizes and sequence lengths.
+We report latency (s/100 tokens, lower is better) and throughput (tokens/s, higher is better).
+In the header, ’k’ denotes 1024 tokens (e.g., 4K = 4096).
+Appendix D Datasets Used in Experiments
+In this section, we provide a comprehensive overview of all the tasks and datasets utilized in the experiments in this paper.
+Math & Science
+This task evaluates the model’s ability to tackle mathematical and scientific problems.
+By directly inputting questions and comparing the model’s output with the correct answers, we calculate the model’s Accuracy on these datasets: GSM8K is a dataset for evaluating model’s math-solving skills, featuring 8,000 elementary-level math word problems requiring basic arithmetic and reasoning.
+GPQA tests model’s understanding of physics concepts and problem-solving across various topics, assessing scientific reasoning abilities.
+TheoremQA evaluates model’s grasp and application of mathematical theorems, ranging from simple applications to complex proofs, testing advanced math skills.
+Commonsense Reasoning (CR)
+This task evaluates model’s ability to make deductions and understand everyday situations using implicit knowledge and logical inference.
+TruthfulQA (ThQA) evaluates model’s ability to generate accurate and truthful responses, testing models on distinguishing fact from fiction, especially in areas prone to misconceptions.
+We use BLEU as the metric.
+CoQA assesses model’s ability to understand and respond to questions in a conversational context, focusing on maintaining coherence and context throughout a dialogue.
+We use F1 Score as the metric.
+Single Document QA (Single-Doc QA)
+This task assesses the model’s reading comprehension skills when dealing with a single, extended document.
+NarrativeQA DBLP:journals/tacl/KociskySBDHMG18 is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes.
+Qasper DBLP:conf/naacl/DasigiLBCSG21 is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information.
+We employ F1 Score as the metric for above two datasets.
+Multi-Document QA (Multi-Doc QA)
+This task evaluates the model’s reading comprehension capabilities across multiple extended documents.
+2WikiMultiHopQA (2WKMQA) DBLP:conf/coling/HoNSA20 is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles.
+MuSiQue DBLP:journals/tacl/TrivediBKS22 evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately.
+We leverage F1 Score as the metric for above two datasets.
+Summarization
+This task examines the model’s ability to comprehend and summarize lengthy documents.
+QMSum DBLP:conf/naacl/ZhongYYZMJACLQR21 is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions.
+Multi-News (M-News) DBLP:conf/acl/FabbriLSLR19 is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic.
+We use Rouge-L as the metric for above two datasets.
+Few-Shot Learning (FSL)
+This task assesses the model’s few-shot learning capabilities.
+TriviaQA DBLP:conf/acl/JoshiCWZ17 is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall.
+We use F1 Score as the metric.
+Code
+This task evaluates the model’s ability to complete and generate code.
+LCC DBLP:conf/icml/GuoXD0M23 is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures.
+We use Edit Sim as the metric.
+Appendix E Ablation
 In this section, we present additional ablation study results for Section 3.4.
@@
 As shown in Table 9, setting $k=1$ not only conserves pruning time but also achieves better model performance with a reduced budget.
-Appendix C Freeze the first two layers of LLM
-In this section, we demonstrate the necessity of using DBudgetKV to compress the KV cache starting from the 2nd layer of LLMs.
-We apply DBudgetKV on Llama2-7B-Chat and conduct experiments on GSM8K and CoQA.
-We first perform a case study, followed by a comparison of the effects of not freezing layers of the model on its performance.
-As shown in Table 9, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence.
-However, the budget exceeds 90% in both cases.
-By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 0 and 1 are relatively low, while the budgets from layer 2 to layer 30 are the same and very high, with the last layer being low again.
-We analyze that since the model, in the first two layers, is not yet able to identify truly important tokens, the attention distribution is relatively uniform.
-This may lead to the model discarding truly important tokens during eviction, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
-Table 8: Performance on GSM8K and CoQA using DBudgetKV with different frozen layers.
-Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
-GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
-Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
-CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
-Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
-Based on the above case study, we attempt to freeze certain layers of the model and explore the optimal freezing configuration.
-We continue to test the results of freezing different layers.
-The resuls are shown in Table 8.
-From the table, we can observe that freezing the first two layers achieves a balance between model performance and budget.
-Moreover, freezing the 31st layer does not significantly enhance the model’s performance and instead leads to an increase in budget.
-Consequently, DBudgetKV ultimately opts to begin KV cache compression from the 2nd layer of the model.
-Appendix D Full Algorithm of DBudgetKV
-  Input: Prompt, Threshold $t$
-  Output: Compressed KV Cache
-  Create Empty List $K_{c},V_{c}$
-  for Transformer Layer $L_{i}$ in LLM do
-     $Q^{i},K^{i},V^{i}\leftarrow L_{i}(\text{Prompt})$
-     $R^{i}\leftarrow$ Postion-Based Importance Rank
-     $A_{last}^{i}\leftarrow\operatorname{Attention}(Q^{i}[\dots,-1,:],K^{iT})$
-     $F_{b}^{i}\leftarrow\operatorname{Frobenius}(A_{last}^{i})$
-     $A_{last}^{i}\leftarrow\operatorname{Square}(A_{last}^{i})$
-     Reorder $A_{last}^{i}$ by Rank $R$
-     $A_{cumsum}^{i}\leftarrow\operatorname{Cumsum}(A_{last}^{i})$
-     $A_{cumsum}^{i}\leftarrow\operatorname{Sqrt}(A_{cumsum}^{i})$
-     $A_{ratio}^{i}\leftarrow(F_{b}^{i}-A_{cumsum}^{i})/F_{b}^{i}$
-     $\text{Index}^{i}\leftarrow\operatorname{Max}(\operatorname{Where}(A_{ratio}^{i}<=t))$
-     $K_{c}^{i}\leftarrow\operatorname{Compress}K^{i}\text{by}R^{i}[I:]$
-     $V_{c}^{i}\leftarrow\operatorname{Compress}V^{i}\text{by}R^{i}[I:]$
-     Append $K_{c}^{i},V_{c}^{i}$ to $K_{c},V_{c}$
-  end for
-  return  $K_{c},V_{c}$
-Algorithm 1 DBudgetKV
-Table 9: Case Study.
-GSM8K
-Input
-A robe takes 2 bolts of blue fiber a-
-nd half that much white fiber. How
-many bolts in total does it take?
-Budget
-Layer 0: 67.71
-Layer 1: 82.29
-Layer 2$\sim$30: 95.83
-Layer 31: 37.50
-Avg.: 93.90
-Output
-I have determined by answering the
-answer to the format of bolts bolts b-
-olts…(repeat)
-Ground-Truth 3
-CoQA
-Input
-You are given a story and a question.
-Answer the question as concisely as
-you can…Question: What color was
-Cotton?
-Budget
-Layer 0: 42.14
-Layer 1: 42.54
-Layer 2$\sim$30: 99.19
-Layer 31: 79.64
-Avg.: 95.03
-Output
-Question: Question: What is the que-
-stion: What is the question:
-Groud-Truth White
-Appendix E Limitations
-The main limitation of DBudgetKV stems from the gap between its current compression budget and the true optimal budget.
-This can be seen in Table 2 that for certain scenarios, e.g.
-QMSum with Mistral-7B, DBudgetKV reaches 84.3% budget while 50% budget is also viable with no performance degradation.
-We regard this gap as room for improvement in future work, facilitating more aggressive KV cache compression while ensuring the full-cache performance.
-Another limitation of DBudgetKV is that though it demonstrates almost lossless compression, we rely on empirical experiments and there is no hard guarantee on the full-performance constraint.
-In Table 2, while both Llama3-8B and Qwen2.5-7B even surpass the full-cache performance, DBudgetKV with Mistral-7B shows trivial degradation of 1.5%.
-More robust methods could be developed as future work to further strengthen this constraint.
-Appendix F Broader Impacts
-Our work has made contributions on both a new objective in this line of research, as well as a new method that successfully fulfills this objective.
-Since all current KV cache pruning methods require a pre-defined KV budget threshold, they either have inconsistent and unstable performance across different input lengths, domains and tasks, or they need careful tuning of their budgets per dataset to ensure the full-cache performance.
-Our proposed method evaluated on 13 diverse datasets, including short and long-context QA, coding and math, demonstrates its effectiveness on addressing this limitation, which automatically adjusts the budget to reach full-cache performance regardless of the inputs.
+Appendix F Analysis of Latency and Throughput
+We perform following experiments for a comprehensive analysis of the latency and throughput.
+We follow the setup in FastGen using inputs of NarrativeQA, and conduct end-to-end latency comparison experiments on a single 80GB A100 GPU, using the standard HuggingFace (HF) Accelerate library as the baseline.
+The results are shown in Table 10.
+GSM8K GPQA CoQA NrtvQA QMSum TriviaQA
+Full 75.28 29.02 52.74 24.06 22.27 70.26
+Oursk=1 76.50 30.13 52.86 23.44 21.95 64.24
+Budget 93.2% 86.7% 81.5% 48.7% 15.0% 41.0%
+SLM0.9 72.78 28.79 52.83 23.75 22.48 69.97
+SLMF0.9 76.42 29.24 52.89 22.95 22.47 69.10
+SLM0.5 3.41 6.03 44.68 20.94 21.80 67.79
+SLMF0.5 4.55 3.13 44.59 21.31 21.88 68.04
+SLM0.2 3.21 3.35 40.11 20.12 21.11 66.17
+SLMF0.2 1.21 1.12 38.44 19.52 20.61 65.69
+Table 11: Performance comparison between ReFreeKV and StreamingLLM with the first two layers frozen (SLMF) under difference KV cache budgets ($0.9$, $0.5$ and $0.2$).
+Methods GSM8K CoQA NQA Musique QMSum
+Llama3-70B-Instruct
+Full 89.69 60.36 27.15 29.31 22.52
+Oursk=1 89.76 60.38 26.94 28.88 22.30
+Budget 93.3% 77.2% 58.7% 60.8% 52.9%
+Qwen2.5-32B-Instruct
+Full 91.36 58.03 24.78 40.04 22.84
+Ours 91.81 57.29 22.65 40.54 22.52
+Budget 91.6% 85.0% 68.2% 74.4% 76.8%
+Qwen2.5-72B-Instruct
+Full 90.22 54.14 24.36 42.13 23.93
+Ours 90.30 54.19 24.10 41.70 23.31
+Budget 90.9% 87.6% 63.4% 65.0% 68.6%
+Table 12: Performance of LLMs with scales of 70B, 34B, and 72B across five datasets using ReFreeKV.
+From the table, we can observe that ReFreeKV consistently outperforms the baseline in both latency and throughput.
+Moreover, ReFreeKV holds its advantage when the batch size increases, and its throughput remains robust.
+It is noteworthy that Table 4 in Section 4.2 also demonstrates the trivial overhead of ReFreeKV due to its simple matrix computation during its two-stage pruning process.
+Appendix G Comparison with StreamingLLM
+We perform additional experiments on six datasets using StreamingLLM (with Llama3-8B) that has the first two layers frozen, as an ablation study on the effects of the proposed pruning and halting process.
+As shown in Table 11, despite having the first two layers frozen, StreamingLLM still falls short of achieving the objective proposed in this work, not able to remain optimal consistently across datasets like GSM8K and GPQA as the budget varies.
+It is demonstrated that solely leveraging the position bias as in StreamingLLM is not able to maintain the full performance; the proposed pruning and halting mechanism that integrates the attention norm metric is necessary in achieving the goal for lossless pruning.
+Appendix H Related Work
+KV Cache Compression
+To mitigate the large memory footprint of the KV cache during LLM inference, a prominent line of work has focused on pruning, which selectively discards less important tokens.
+Methods like Scissorhands DBLP:conf/nips/LiuDLWXXKS23 and SnapKV DBLP:conf/nips/LiHYVLYCLC24 retain tokens based on high attention scores.
+Others employ strategies based on token recency and historical importance, such as H2O DBLP:conf/nips/Zhang00CZC0TRBW23 or StreamingLLM DBLP:conf/iclr/XiaoTCHL24.
+FastGen DBLP:conf/iclr/Ge0LZ0024 further refines this by adapting retention strategies on a per-head basis.
+A common limitation of these pruning methods is their reliance on a pre-defined memory budget, which can lead to inconsistent performance across different domains and datasets.
+In contrast, our work lifts this constraint, aiming for near full-cache performance with a dynamically sized budget without requiring task-specific tuning.
+Beyond pruning, other compression paradigms have been explored.
+These include quantization-based methods, which reduce the numerical precision of cached values DBLP:journals/corr/abs-2405-06219; DBLP:conf/nips/HooperKMMSKG24; DBLP:journals/corr/abs-2503-18773; DBLP:journals/corr/abs-2402-12065; DBLP:journals/corr/abs-2502-04420, and low-rank approximation techniques that apply matrix decomposition to compress the KV cache DBLP:conf/iclr/ChangLLCHWHCAW25; DBLP:journals/corr/abs-2405-04434.
+Attention Patterns
+Our approach is grounded in research on the structural patterns of attention mechanisms.
+Studies have shown that attention heads often specialize, with some exhibiting predictable, position-biased behaviors such as consistently focusing on adjacent tokens DBLP:journals/corr/abs-1911-02150.
+This indicates a degree of redundancy in their learned patterns.
+Furthermore, the importance of individual heads can vary significantly across tasks DBLP:journals/corr/abs-2404-07965, suggesting opportunities for dynamic optimization.